### v2

# Injectivity of ReLU-layers: Perspectives from Frame Theory

Abstract
Injectivity is the defining property of a mapping that ensures no information is lost and any input can be perfectly reconstructed from its output. By performing hard thresholding, the ReLU function naturally interferes with this property, making the injectivity analysis of ReLU-layers in neural networks a challenging yet intriguing task that has not yet been fully solved. This article establishes a frame theoretic perspective to approach this problem. The main objective is to develop the most general characterization of the injectivity behavior of ReLU-layers in terms of all three involved ingredients: $(i)$ the weights, $(ii)$ the bias, and $(iii)$ the domain where the data is drawn from. Maintaining a focus on practical applications, we limit our attention to bounded domains and present two methods for numerically approximating a maximal bias for given weights and data domains. These methods provide sufficient conditions for the injectivity of a ReLU-layer on those domains and yield a novel practical methodology for studying the information loss in ReLU layers. Finally, we derive explicit reconstruction formulas based on the duality concept from frame theory.
1 Introduction
The Rectified Linear Unit defined as $\text{ReLU}(s)=\max(0,s)$ for $s\in\mathbb{R}$ in the role as a non-linear activation function in neural networks has become indispensable in modern deep learning [14, 20, 15, 24]. Since originally introduced as a way to regularize the gradients in deep network architectures, there have been hardly any networks that do not use ReLU activation or some derivation of it [12, 17, 22].
A ReLU-layer $C_{\alpha}(x)=\text{ReLU}(Cx-\alpha)$ is the composition of an affine linear map comprising the multiplication by a weight matrix $C\in\mathbb{R}^{m\times n}$ and the shift by a bias vector $\alpha\in\mathbb{R}^{m}$, and an entry-wise application of ReLU on its output. The injectivity of such a ReLU-layer, and with that, the possibility of inverting it and inferring $x$ from $C_{\alpha}(x)$, is a desired property in various applications. As such, there has been interest in building injective models and inverting them on their range to regularize ill-posed inverse problems or designing injective generative models, such as normalizing flows, for manifold learning or compressed sensing [19, 26, 5]. Generally, knowing if a layer involves a loss of information or not increases the interpretability of the network immensely. As such, one can use an inverse mapping to trace back each layer output to its source input, which can help to decipher the decision-making process, diagnose model behavior, identify biases, and study accountability. A full solution, especially one for verifying injectivity of a given ReLU-layer in practice, has not been found so far.
The goal of this paper is to demystify the injectivity of a single ReLU-layer with any given weights, biases, and inputs from bounded domains in the most general way. This involves a thorough analysis of the fundamental properties with respect to applications, and the derivation of different characterizations with regard to all properties involved, namely weights, biases and input domain. Wherever possible, the theoretical results presented are translated into applications by means of algorithmic solutions, yielding novel ways of verifying injectivity on bounded input domains in practice. Explicit reconstruction formulas and their implementation, together with a stability analysis for injective ReLU-layers complete the claim of the paper.
The methodology to achieve the results is based on frame theory, a mathematical paradigm that deals with stable and invertible representations in Hilbert spaces [11]. We start to motivate the approach from the matrix perspective. When the weight matrix $C$ is a “thin” matrix, meaning it has more rows than columns ($m\geq n$) and has full rank, its row vectors form a spanning set for the domain space $\mathbb{R}^{n}$. Equivalently, the associated linear transform is injective. In the context of frame theory, we say that these vectors form a frame for $\mathbb{R}^{n}$ [11], and $C$ is referred to as the associated analysis operator. This operator maps an input $x$ to its frame coefficients as
$Cx=\begin{pmatrix}-\phi_{1}-\\ \vdots\\ -\phi_{m}-\end{pmatrix}x=(\langle x,\phi_{i}\rangle)_{i=1}^{m}.$ (1)
We write $\Phi=(\phi_{i})_{i=1}^{m}$ and interpret the transform above as measuring the correlation of $x$ to all frame elements $\phi_{i}$ via the sequence $(\langle x,\phi_{i}\rangle)_{i=1}^{m}$. These coefficients give a redundant representation of $x$, from which we can always infer $x$ explicitly. The concept of a frame offers the possibility to “quantify” injectivity of a redundant representation in terms of its numerical stability, and how to explicitly reconstruct the input vector from its representations. Hence, it provides exactly the right tools for the goal of the paper. Despite frame theory being a “linear theory”, it has shown to be highly suitable in non-linear settings as well. This has been done extensively for phase-retrieval [3], where one asks for the injectivity of
$\begin{split}C_{M}:\mathbb{R}^{n}/\{\pm 1\}&\rightarrow\mathbb{R}^{m}\\ x&\mapsto\left(\big{|}\langle x,\phi_{i}\rangle\big{|}\right)_{i=1}^{m}.\end{split}$ (2)
It is well known that $C_{M}$ is injective if and only if for any partition of the frame $\Phi$ into two sub-collections, at least one of them still forms a frame [3]. Inspired by this approach, we may write a ReLU-layer as
$\begin{split}C_{\alpha}:\mathbb{R}^{n}&\rightarrow\mathbb{R}^{m}\\ x&\mapsto\left(\text{ReLU}(\langle x,\phi_{i}\rangle-\alpha_{i})\right)_{i=1}^{m}.\end{split}$ (3)
In [25] (using other terminology) it has been shown that $C_{\alpha}$ is injective if and only if for any $x\in\mathbb{R}^{n}$ the frame elements that are not affected by ReLU still form a frame. We will call a frame with this property $\alpha$-rectifying on $\mathbb{R}^{n}$. This property will be the main focus of the paper.
A fundamental assumption that is key in our work is that when working with data sets, it is not necessarily the most informative approach to consider the entire $\mathbb{R}^{n}$ as the domain where a ReLU-layer should be injective. As such, when considering standard normalization schemes for training and testing neural networks, it is more interesting to study injectivity on bounded subsets $K\subseteq\mathbb{R}^{n}$, where we assume the data to lie. We show that the choice of the data domain can have a profound impact on the injectivity behavior of $C_{\alpha}$, e.g., it might not be injective on $\mathbb{R}^{n}$ but on some bounded subset of interest, and how to achieve this. Prototypical examples of such domains include the closed ball of radius $r>0$ in $\mathbb{R}^{n}$, given by $\mathbb{B}_{r}=\{x\in\mathbb{R}^{n}:\|x\|\leq r\}$, the closed donut by excluding small samples $\mathbb{D}_{r,s}=\overline{\mathbb{B}_{r}\setminus\mathbb{B}_{s}}$ with $s<r$, or the sphere $\mathbb{S}=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ [21, 18]. Since the output of a ReLU-layer has only non-negative entries it is also important to consider the non-negative closed ball $\mathbb{B}_{r}^{+}=\mathbb{B}_{r}\cap\mathbb{R}^{n}_{+}$ as the input domain of a ReLU-layer, which applies to the output of the first one. The restriction of the domain of $C_{\alpha}$ from $\mathbb{R}^{n}$ to a bounded $K\subseteq\mathbb{R}^{n}$ increases the feasibility and applicability of the problem in practice, while also making the mathematical setting more versatile. Furthermore, it establishes a natural connection to the bias vector and provides a framework where we can control the injectivity behavior through these two ingredients. This in turn allows us to approach the injectivity analysis also algorithmically, hence, can be implemented and used in practical applications.
The main theoretical component of this paper is the different characterizations of the injectivity of a ReLU-layer as deterministic operator, summarized in the following theorem.
Theorem 1.
Let $\Phi=(\phi_{i})_{i\in I}\subset\mathbb{R}^{n}$ have a unique most correlated basis everywhere (see Definition 3), $\emptyset\neq K\subseteq\mathbb{R}^{n}$ be open or strictly convex, $K_{i}^{\sharp}\neq\emptyset$ for all $i\in I$ (see Equations (30) (32)), and $\alpha\in\mathbb{R}^{m}$. The following are equivalent.
  1. (i)
    The ReLU-layer $C_{\alpha}$ associated with the frame $\Phi$ and the bias $\alpha$ is injective on the domain $K$.
  2. (ii)
    The frame $\Phi$ is $\alpha$-rectifying on $K$, i.e., for every $x\in K$, the active frame elements $\Phi_{I_{x}^{\alpha}}$ form a frame (Definition 2, Theorem 3, 2).
  3. (iii)
    The domain $K$ lies in the maximal domain $\mathcal{K}_{\alpha}^{*}$ (Theorem 4).
  4. (iv)
    The values of the bias $\alpha$ are smaller or equal to the values of the maximal bias $\alpha_{K}^{\sharp}$ (Theorem 5).
Moreover, for any bias $\alpha$ the maximal domain $\mathcal{K}_{\alpha}^{*}$ can be constructed explicitly as the union of intersections of closed affine half-spaces, and for any domain $K$ the maximal bias $\alpha_{K}^{\sharp}$ can be approximated numerically via sampling or via the inscribing polytope associated with $\Phi$.
Using the duality concept from frame theory we derive inversion formulas, or ReLU-duals, for injective ReLU-layers that can be implemented easily. The main practical component of the paper is the numerical computation of a bias $\alpha_{K}^{\flat}$ which is close to $\alpha_{K}^{\sharp}$ from the theorem above but of a simpler form. This allows for the study of the injectivity behavior of a ReLU-layer in practice, and in turn, the effective control thereof.
Related Work
The approach in classical phase-retrieval in $\mathbb{R}^{n}$ by Balan et al. in [3] was decisive for the idea of characterizing the injectivity of a ReLU-layer in terms of a property of the associated frame. The same approach is taken by Alharbi et al. to study the recovery of vectors from saturated inner measurements [2]. In a machine learning context, Puthawala et al. have introduced the notion of directed spanning sets in [25] as an equivalent concept to the admissibility condition by Bruna et al. in [7] to characterize a ReLU-layer to be injective on $\mathbb{R}^{n}$. While the primary goal in [25] is the study of globally injective ReLU-networks on $\mathbb{R}^{n}$, and the one in [7] is a Lipschitz stability analysis of ReLU-layers, our goal is to demystify the injectivity of a single ReLU-layer in a more realistic setting, namely with any given weights and biases on bounded input data domains $K\subseteq\mathbb{R}^{n}$, and to provide methods of verifying injectivity in practice. Further related preprints are by Behrmann et al., who study the pre-images of ReLU-layers from a geometric point of view [4] and by Maillard et al., which focus on injectivity of ReLU-layers with random weights [23].
Outline
The paper is divided into four sections. In Section 2, we introduce $\alpha$-rectifying frames as a characterizing family of frames that are associated with ReLU-layers that are injective on $K\subseteq\mathbb{R}^{n}$ and discuss fundamental properties. In Section 3 we study the interplay of the domain $K$ and the bias vector $\alpha$ and provide statements about the maximal domain and the maximal bias such that the associated ReLU-layers become critically injective. We present two methods to approximate the maximal bias and provide algorithmic solutions to apply them in practice. Section 4 is dedicated to explicit reconstruction formulas for injective ReLU-layers and a stability analysis of the recovery map.
2 Frames and the Injectivity of ReLU-layers
We can reconstruct any input $x\in\mathbb{R}^{n}$ after the multiplication by a matrix $C\in\mathbb{R}^{m\times n}$ if and only if the collection of column vectors of $C$ spans $\mathbb{R}^{n}$. In other words, $C$ is injective. Frame theory offers a definition of a spanning set that allows us to quantify “how good” it spans $\mathbb{R}^{n}$. Throughout this paper, we denote a finite collection of $m$ vectors as $\Phi=(\phi_{i})_{i\in I}\subset\mathbb{R}^{n}$, using index sets $I$ with $|I|=m\geq n$. Then, $\Phi$ forms a frame for $\mathbb{R}^{n}$ if there are constants $0<A\leq B$, called the frame bounds of $\Phi$, such that
$A\cdot\|x\|^{2}\leq\sum_{i\in I}|\langle x,\phi_{i}\rangle|^{2}\leq B\cdot\|x\|^{2}$ (4)
holds for all $x\in\mathbb{R}^{n}$ [10]. It is easy to see that (4) is equivalent to $\Phi$ being a spanning set. Moreover, we note that the frame bounds $A,B$ reflect the numerical stability properties of the representation of $x$ under $\Phi$. For $J\subseteq I$, we denote by $\Phi_{J}=(\phi_{i})_{i\in J}$ the sub-collection of $\Phi$ with respect to the index set $J$. Any frame with $m=n$ is a basis and we will always assume that $0\notin\Phi$. The mapping that maps a vector $x$ to its frame coefficients $\left(\langle x,\phi_{i}\rangle\right)_{i\in I}$ is called the analysis operator associated with $\Phi$. In our setting, we can identify the operator application on $x$ with the multiplication of $x$ by the weight matrix $C$ from the left, resulting in
$\begin{split}C:\mathbb{R}^{n}&\rightarrow\mathbb{R}^{m}\\ x&\mapsto\left(\langle x,\phi_{i}\rangle\right)_{i\in I}.\end{split}$ (5)
In this wording, $\Phi$ is a frame, if and only if $C$ is injective. This has motivated us to define a ReLU-layer as follows.
Definition 1 (ReLU-layer).
A ReLU-layer associated with a collection of vectors $\Phi=(\phi_{i})_{i\in I}\subset\mathbb{R}^{n}$, a bias vector $\alpha=(\alpha_{1},...,\alpha_{m})^{\top}\in\mathbb{R}^{m}$ and an input domain $K\subseteq\mathbb{R}^{n}$ is defined as the non-linear operator given by
$C_{\alpha}:K$ $\rightarrow\mathbb{R}^{m}$
$x$ $\mapsto\left(\max(0,\langle x,\phi_{i}\rangle-\alpha_{i})\right)_{i\in I}.$
To encode the injectivity of $C_{\alpha}$ directly in terms of $\Phi$ we introduce a customized family of frames called $\alpha$-rectifying frames.
2.1 Alpha-rectifying Frames
The shift by the bias $\alpha$ and the application of the ReLU function on the frame analysis coefficients act as a thresholding mechanism that neglects all frame elements $\phi_{i}$ where $\langle x,\phi_{i}\rangle<\alpha_{i}$ for given $x\in K$. According to this observation, for $x\in\mathbb{R}^{n}$ and $\alpha\in\mathbb{R}^{m}$ we denote the index set associated with the frame elements that are active for $x$ and $\alpha$ by
$I_{x}^{\alpha}=\{i\in I:\langle x,\phi_{i}\rangle\geq\alpha_{i}\}.$ (6)
Dual to this notion, for $i\in I$ and $\alpha\in\mathbb{R}^{m}$ we denote the closed affine half-space of points where the frame element $\phi_{i}$ is active by
$\Omega_{i}^{\alpha}=\{x\in\mathbb{R}^{n}:\langle x,\phi_{i}\rangle\geq\alpha_{i}\}.$ (7)
Definition 2 ($\alpha$-rectifying frames).
The collection $\Phi=(\phi_{i})_{i\in I}\subset\mathbb{R}^{n}$ is called $\alpha$-rectifying on $K\subseteq\mathbb{R}^{n}$ for $\alpha\in\mathbb{R}^{m}$ if for all $x\in K$ the sub-collection of active frame elements $\Phi_{I_{x}^{\alpha}}=(\phi_{i})_{i\in I_{x}^{\alpha}}$ forms a frame for $\mathbb{R}^{n}$.
This definition gives an alternative perspective to [25, Definition 1]. We note that if $K$ lies in a proper subspace of $\mathbb{R}^{n}$, then it is sufficient that $\Phi_{I_{x}^{\alpha}}$ is a frame for that subspace for every $x\in K$. Furthermore, if $\Phi$ is $\alpha$-rectifying on $K$ then $\Phi$ is $\alpha^{\prime}$-rectifying on $K$ for all $\alpha^{\prime}$ with $\alpha^{\prime}\leq\alpha$. Hence, we are naturally interested in knowing the largest bias that allows this property. Figure 1 illustrates the two notions in (6) and (7) on the closed unit ball $\mathbb{B}$ in $\mathbb{R}^{2}$, and a simple example of an $\alpha$-rectifying frame.
Figure 1: Illustrations of the notions $I_{x}^{\alpha}$ and $\Omega_{i}^{\alpha}$ related to active frame elements on $\mathbb{B}$. The frame $(\phi_{1},\phi_{2})$ in the most right example is $\alpha$-rectifying on $K$ if $K\subseteq(\Omega_{1}^{\alpha}\cap\Omega_{2}^{\alpha})$.
We now present the fundamental connection between the $\alpha$-rectifying property and the injectivity of $C_{\alpha}$ on $K$, forming the backbone of this paper. Theorems 3 and 2 address the two directions of this implication separately, each focusing on specific properties of $K$, thereby generalizing [25, Theorem 2].
Theorem 2 (Injectivity of ReLU-layers I).
Given $\Phi=(\phi_{i})_{i\in I}\subset\mathbb{R}^{n}$, $\alpha\in\mathbb{R}^{m}$, and $\emptyset\neq K\subseteq\mathbb{R}^{n}$. If $\Phi$ is $\alpha$-rectifying on $K$, then $C_{\beta}$ is injective on $K$ for all $\beta<\alpha$ (i.e., $\beta_{i}<\alpha_{i}$ for all $i\in I$). Moreover, if $K$ is open or convex, $C_{\alpha}$ is injective on $K$.
[…]
Theorem 3 (Injectivity of ReLU-layers II).
Given $\Phi=(\phi_{i})_{i\in I}\subset\mathbb{R}^{n}$, $\alpha\in\mathbb{R}^{m}$, and let $\emptyset\neq K\subseteq\mathbb{R}^{n}$, $|K|\geq 2$ be open or strictly convex. If $C_{\alpha}$ is injective on $K$, then $\Phi$ is $\alpha$-rectifying on $K$.
Theorem 2.
Assume $C_{\beta}x=C_{\beta}y$, for $x,y\in K$ and let $\alpha>\beta$. By definition, for all $i\in I^{\alpha}_{x}$ it holds that $\langle x,\phi_{i}\rangle\geq\alpha_{i}>\beta_{i}.$ From the fact that
$\langle x,\phi_{i}\rangle>\beta_{i}\;\text{ if and only if }\;\langle y,\phi_{i}\rangle>\beta_{i},$ (8)
we deduce that $\langle x,\phi_{i}\rangle=\langle y,\phi_{i}\rangle$ for all $i\in I^{\beta}_{x}$. Assuming that $\Phi$ is $\alpha$-rectifying, $\Phi_{I^{\alpha}_{x}}$ is a frame, hence so is $\Phi_{I^{\beta}_{x}}$. It follows that $x=y$. Therefore, $C_{\beta}$ is injective on $K$.
To show the moreover part, let $C_{\alpha}x=C_{\alpha}y$, for $x,y\in K$. Clearly,
$\langle x,\phi_{i}\rangle=\langle y,\phi_{i}\rangle$ (9)
for all $i\in I_{x}^{\alpha}\cap I_{y}^{\alpha}$. We will show that if $K$ is open or convex, then $\Phi_{I_{x}^{\alpha}\cap I_{y}^{\alpha}}$ is a frame.
So let us first assume that $K$ is open. Then we may choose $\varepsilon>0$ such that the open ball around $x$, denoted by $B_{\varepsilon}^{\circ}(x)$ is contained in $K$. By contradiction, we assume $x\neq y$. Let $0<\delta<\varepsilon\cdot\|x-y\|^{-1}$, then
$x_{\delta}:=(1-\delta)x+\delta y\in B_{\varepsilon}^{\circ}(x).$ (10)
Now let $i\in I_{x_{\delta}}^{\alpha}$. By the linearity of the inner product, we have the following.
If $\text{If }\;\langle x_{\delta},\phi_{i}\rangle$ $>\alpha_{i}\;\text{ then }\;\langle x,\phi_{i}\rangle>\alpha_{i}\text{ and }\langle y,\phi_{i}\rangle>\alpha_{i}.$ (11)
If $\text{If }\;\langle x_{\delta},\phi_{i}\rangle$ $=\alpha_{i}\;\text{ then }\;\langle x,\phi_{i}\rangle=\alpha_{i}\text{ and }\langle y,\phi_{i}\rangle=\alpha_{i}.$ (12)
Therefore, $I_{x_{\delta}}\subseteq I_{x}^{\alpha}\cap I_{y}^{\alpha}$. By (10), $x_{\delta}\in K$ and $\Phi_{I_{x_{\delta}}}$ is a frame, hence, so is $\Phi_{I_{x}^{\alpha}\cap I_{y}^{\alpha}}$.
Now let us assume $K$ to be convex. For $\lambda\in(0,1)$,
$x_{\lambda}:=(1-\lambda)x+\lambda y\in K.$ (13)
By the same arguments as above, (11) and (12) hold for $\langle x_{\lambda},\phi_{i}\rangle$, hence $I_{x_{\lambda}}\subseteq I_{x}^{\alpha}\cap I_{y}^{\alpha}$. By (13), $\Phi_{I_{x}^{\alpha}\cap I_{y}^{\alpha}}$ is a frame. So for both cases, we can deduce that $x=y$, hence $C_{\alpha}$ is injective. ∎
Theorem 3.
We prove the claim by contraposition. Assume that $\Phi$ is not $\alpha$-rectifying. Then there is $x\in K$ such that $(\phi_{i})_{i\in I_{x}^{\alpha}}$ is not a frame. Hence, there is
$0\neq r\in\text{span}(\phi_{i})_{i\in I_{x}^{\alpha}}^{\bot}.$ (14)
If $K$ is open, for all sufficiently small $\varepsilon>0$ we have that
$y_{\pm}:=x\pm\varepsilon r\in K.$
For $i\in I_{x}^{\alpha}$, (14) implies that $\langle y_{+},\phi_{i}\rangle=\langle y_{-},\phi_{i}\rangle,$ leading to
$\max(0,\langle y_{+},\phi_{i}\rangle-\alpha_{i})=\langle x,\phi_{i}\rangle-\alpha_{i}=\max(0,\langle y_{-},\phi_{i}\rangle-\alpha_{i}).$ (15)
If $I_{x}^{\alpha}=I$, then (15) already implies $C_{\alpha}y_{+}=C_{\alpha}y_{-}$, so that $C_{\alpha}$ is not injective on $K$.
To address the case $I_{x}^{\alpha}\subset I$, we recall that $\langle x,\phi_{i}\rangle<\alpha_{i}$ holds for all $i\in I\setminus I_{x}^{\alpha}$. Therefore, we may choose $\varepsilon$ sufficiently small, such that $y_{\pm}\in K$ and
$\langle y_{\pm},\phi_{i}\rangle<\alpha_{i},\qquad i\in I\setminus I_{x}^{\alpha}.$ (16)
Observation (16) leads to
$0=\max(0,\langle y_{\pm},\phi_{i}\rangle-\alpha_{i}),\qquad i\in I\setminus I_{x}^{\alpha}.$ (17)
According to (15) and (17), we derive $C_{\alpha}y_{+}=C_{\alpha}y_{-}$. Thus, $C_{\alpha}$ is not injective on $K$.
If $K$ is strictly convex and contains more than one element, then for every $z\in K$ where $z\neq x$ and $\lambda\in(0,1)$ with the same assumptions on $x$ as above,
$x_{\lambda}:=(1-\lambda)x-\lambda z\in\mathring{K}.$
Using the linearity of the inner product, there is $\lambda$ sufficiently small such that
$\langle x_{\lambda},\phi_{i}\rangle<\alpha_{i},\qquad i\in I\setminus I_{x}^{\alpha}.$ (18)
Since $\mathring{K}$ is open and not empty, we can apply the same argument as above with $x_{\lambda}$ to show that $C_{\alpha}$ is not injective on $K$. ∎
This shows that the injectivity of a ReLU-layer is contingent upon topological properties of the domain $K$ from which the data is drawn. For the cases $K=\mathbb{R}^{n}$ and $\mathbb{B}_{r}$, we have established equivalence, where the former case corresponds to [25, Theorem 2]. Since the non-negative ball $\mathbb{B}_{r}^{+}=\mathbb{B}_{r}\cap\mathbb{R}^{n}_{+}$ is not open or strictly convex, only one direction holds. In the case of the donut $\mathbb{D}_{r,s}=\overline{\mathbb{B}_{r}\setminus\mathbb{B}_{s}}$, which is not open and not convex, only the first part of Theorem 2 holds.
While the $\alpha$-rectifying property makes the injectivity of a ReLU-layer more accessible, it remains challenging to verify it in practice. In the following, we discuss how specific properties of the underlying frame influence its $\alpha$-rectifying property, and how to make use of these properties.
2.2 Normalized Frames
It is a simple, yet, crucial observation that we may restrict the $\alpha$-rectifying property to frames with unit norm vectors by scaling the bias with the norms of the frame elements. In the context of neural networks, this corresponds to a normalization of the weight matrix, which is a standard normalization technique [29].
Lemma 1.
A collection $\Phi=(\phi_{i})_{i\in I}$ is $\alpha$-rectifying on $K$ if and only if the normalized collection $\Phi^{\prime}=\left(\phi_{i}\cdot\|\phi_{i}\|^{-1}\right)_{i\in I}$ is $\alpha^{\prime}$-rectifying on $K$, where $\alpha^{\prime}_{i}=\alpha_{i}\cdot\|\phi_{i}\|^{-1}$.
The statement follows from the fact that $\langle x,\phi_{i}\rangle\geq\alpha_{i}$ is equivalent to $\langle x,\phi_{i}\cdot\|\phi_{i}\|^{-1}\rangle\geq\alpha_{i}\cdot\|\phi_{i}\|^{-1}$ for all $x\in K$. Therefore, we may always assume $\Phi\subset\mathbb{S}$, i.e., $\|\phi_{i}\|=1$ for all $i\in I$, which simplifies the problem setting substantially. If needed, the norms can be reintroduced at any stage of processing or analysis.
2.3 Full-spark Frames
The effect of ReLU can be interpreted as introducing input-dependent erasures in the underlying frame. The smallest number $s\geq n$ such that any sub-collection of $\Phi$ with $s$ elements is still a frame for $\mathbb{R}^{n}$ is called the spark of $\Phi$. Frames with $s=n$ are called full-spark frames. This family of frames has shown to be maximally robust to erasures [1] which makes the full-spark property interesting for injective ReLU-layers in particular. If a full-spark frame has $m\geq 2n-1$ elements then the phase-retrieval operator (2) is injective [3]. For the ReLU case, knowing the spark of $\Phi$ relaxes the condition for a frame to be $\alpha$-rectifying to a counting argument.
Corollary 1.
Let $\Phi$ be a frame with spark $s$ then $\Phi$ is $\alpha$-rectifying on $K$ if and only if $|I_{x}^{\alpha}|\geq s$ for all $x\in K$.
Although it is an NP-hard problem to verify if a given frame is full-spark, in a numerical setting it is a mild condition that is almost surely satisfied in the presence of randomness and quantization. As such, if the entries of the frame elements are i.i.d. samples from an absolutely continuous probability distribution, then the associated random frame is full-spark with probability one [1]. Since most initialization methods in neural networks are based on i.i.d. sampling schemes, full-spark frames appear naturally in the context of deep learning.
2.4 Perturbed Frames
For a bounded domain $K$, small perturbations of an $\alpha$-rectifying frame result in an $\alpha^{\prime}$-rectifying frame where $\alpha^{\prime}$ is close to $\alpha$. In other words, the $\alpha$-rectifying property is robust to perturbation.
Lemma 2.
Let $\Phi=(\phi_{i})_{i\in I}$ be $\alpha$-rectifying on a bounded domain $K$ with $M=\sup_{x\in K}\|x\|$. For $\varepsilon>0$, a perturbed frame $\Phi^{\prime}=(\phi^{\prime}_{i})_{i\in I}$ satisfying $\|\phi_{i}-\phi^{\prime}_{i}\|<\varepsilon$ for all $i\in I$ is $\alpha^{\prime}$-rectifying on $K$ with $\alpha^{\prime}_{i}=\alpha_{i}-\varepsilon\cdot M,\;i\in I$.
Proof.
Let $x\in K$, then for any $i\in I_{x}^{\alpha}$ it holds that
$\langle x,\phi^{\prime}_{i}\rangle=\langle x,\phi_{i}\rangle-\langle x,\phi_{i}-\phi^{\prime}_{i}\rangle>\alpha_{i}-\varepsilon\cdot\|x\|\geq\alpha_{i}-\varepsilon\cdot M.$
It is known that for any frame $\Phi$ there is an arbitrarily small perturbation such that the resulting perturbed frame is full-spark [10]. Therefore, we may interpret Lemma 2 in the sense that for any $\alpha$-rectifying frame, there is an arbitrarily close full-spark frame that is $\alpha^{\prime}$-rectifying with $\alpha^{\prime}$ arbitrary close to $\alpha$.
2.5 Redundancy
One of the central properties of a frame is its redundancy, i.e., the ratio $q=\frac{m}{n}\geq 1$. For random ReLU-layers with i.i.d. Gaussian entries and no bias, it has been studied at which redundancy they become asymptotically injective on $\mathbb{R}^{n}$ (for $n\rightarrow\infty$) [25, 23]. In particular, let $p_{m,n}$ denote the probability that $C_{\alpha}$ with $m$ frame elements in $\mathbb{R}^{n}$ and $\alpha\equiv 0$ is injective then it has been proven that
$\left(\frac{m}{n}\leq 3.3\Rightarrow\lim_{n\rightarrow\infty}p_{m,n}=0\quad\text{ and }\quad\frac{m}{n}\geq 9.091\Rightarrow\lim_{n\rightarrow\infty}p_{m,n}=1\right).$
Furthermore, the authors in [23] state the conjecture that there exists a redundancy $q\in(6.6979,6.6981)$ where the transition from non-injectivity to injectivity happens (asymptotically). We will revisit this conjecture in the experimental part later in the paper (Section 3.3).
Given any frame $\Phi$ and $\alpha\equiv 0$, a trivial construction of a frame that is $\alpha$-rectifying on $\mathbb{R}^{n}$ is $\Psi=\left(\Phi,-\Phi\right)$ [7, 25, 31]. Doubling the redundancy in this symmetric way makes a ReLU-layer trivially injective by design. However, as the $\alpha$-rectifying property depends heavily on the geometric characteristics of the frame, it is difficult to establish sufficient conditions for the injectivity of $C_{\alpha}$ only in terms of the redundancy, in contrast to the phase-retrieval operator. On the other hand, as already noted in [7], a redundancy of two is necessary for injectivity on $K=\mathbb{R}^{n}$. In the following proposition, we extend this statement from $\alpha\equiv 0$ to arbitrary $\alpha$.
Proposition 1.
Any $\alpha$-rectifying frame on $\mathbb{R}^{n}$ has at least redundancy two.
Proof.
Choose $x\in\mathbb{R}^{n}$ with $\langle x,\phi_{i}\rangle\neq 0$ for all $i\in I$ and let $\alpha\in\mathbb{R}^{m}$. We denote
$I_{x}^{+}$ =$=\{i\in I:\langle x,\phi_{i}\rangle>0\},$ (19)
$I_{x}^{-}$ =$=\{i\in I:\langle x,\phi_{i}\rangle<0\}.$ (20)
By the choice of $x$, the sets $I_{x}^{+}$ and $I_{x}^{-}$ form a disjoint partition of $I$. We define
$t^{*}:=\max_{\begin{subarray}{c}i\in I\end{subarray}}\frac{\alpha_{i}}{\langle x,\phi_{i}\rangle}$ (21)
then for all $t>t^{*}>0$ we have
$\langle tx,\phi_{i}\rangle>\alpha_{i}\;\text{ and }\;\langle-tx,\phi_{i}\rangle<\alpha_{i}\quad\text{for }i\in I_{x}^{+}$ (22)
$\langle tx,\phi_{i}\rangle<\alpha_{i}\;\text{ and }\;\langle-tx,\phi_{i}\rangle>\alpha_{i}\quad\text{for }i\in I_{x}^{-}.$ (23)
Hence, $I_{tx}^{\alpha}=I_{x}^{+}$ and $I_{-tx}^{\alpha}=I_{x}^{-}$. We found two elements $u=tx,v=-tx\in\mathbb{R}^{n}$ with $I_{u}^{\alpha}\cap I_{v}^{\alpha}=\emptyset$. Assuming $\Phi$ to be $\alpha$-rectifying on $\mathbb{R}^{n}$ implies that $\Phi_{I_{u}^{\alpha}}$ and $\Phi_{I_{v}^{\alpha}}$ are frames, i.e., contain at least $n$ elements in particular. Since $I_{u}^{\alpha}\cap I_{v}^{\alpha}=\emptyset$, it must hold that $m\geq 2n$. ∎
Assuming that the input for a ReLU-layer is contained in a closed ball $\mathbb{B}_{r}$, we find that the necessary redundancy-two condition from Proposition 1 breaks.
Lemma 3.
Any normalized frame is $\alpha$-rectifying on $\mathbb{B}_{r}$ for $\alpha\equiv-r$. If it is a basis, this is also necessary.
Proof.
Let $x\in\mathbb{B}_{r}$. Since $\langle x,\phi_{i}\rangle=\|x\|\langle\frac{x}{\|x\|},\phi_{i}\rangle\geq-\|x\|\geq-r$ the first statement follows. For the second, let $\Phi$ be a basis, then $\Phi$ is $\alpha$-rectifying on $\mathbb{B}_{r}$ if and only if for every $x\in\mathbb{B}_{r}$ it holds that $I=I_{x}^{\alpha}$. In particular, since $-r\cdot\Phi\subset\mathbb{B}_{r}$ and for every $i\in I$, the maximal choice of the $\alpha_{i}$ is determined by the fact that $\langle-r\cdot\phi_{i},\phi_{i}\rangle=-r$. ∎
This emphasizes that the choice of the input domain can have a significant impact on the $\alpha$-rectifying property. The following section will examine this interaction between domain and bias in greater detail and explain how it can be exploited to our advantage.
3 Interplay of Domain and Bias
In the context of applications, we may find ourselves in a situation where we have provided a trained ReLU layer and wish to ascertain whether it is injective for a specific set of data points. One way to address this is to verify that the data points in question are contained within a set where we have already established that the ReLU-layer is injective. This leads to the following question.
Q1: Given $\alpha$, what is the largest domain $K$ such that $\Phi$ is $\alpha$-rectifying on $K$?
An alternative approach is to ascertain whether the bias is smaller than one for which we already know that the corresponding ReLU-layer is injective. This leads to the dual question to the one above.
Q2: Given $K$, what is the largest bias $\alpha$ such that $\Phi$ is $\alpha$-rectifying on $K$?
Answering these two questions provides characterizations of the $\alpha$-rectifying property in terms of domain and bias, respectively, yielding alternative ways of verifying the injectivity of the associated ReLU-layer. To better understand how bias and domain interact, we point out some basic scaling relations.
Lemma 4.
Let $\Phi$ be $\alpha$-rectifying on $K$. The following holds.
  1. (i)
    $\Phi$ is $(r\cdot\alpha)$-rectifying on $r\cdot K$ for any $r>0$.
  2. (ii)
    If $\alpha\geq 0$, then $\Phi$ is $\alpha$-rectifying on $r\cdot K$ with $r\geq 1$.
  3. (iii)
    If $0\in K$, then there is a sub-frame $\Phi_{J}\subseteq\Phi$ with $\alpha_{i}\leq 0$ for all $i\in J$.
Proof.
All properties are easy to see.
  1. (i)
    For $r>0$, $\langle x,\phi_{i}\rangle\geq\alpha_{i}\Leftrightarrow\langle r\cdot x,\phi_{i}\rangle\geq r\cdot\alpha_{i}$.
  2. (ii)
    For $r\geq 1$, $\langle x,\phi_{i}\rangle\geq\alpha_{i}\Rightarrow\langle r\cdot x,\phi_{i}\rangle\geq r\cdot\alpha_{i}\geq\alpha_{i}$.
  3. (iii)
    For $x=0$, $\Phi_{I_{x}^{\alpha}}$ is a frame with $0=\langle 0,\phi_{i}\rangle\geq\alpha_{i}$ for all $i\in I_{x}^{\alpha}$.
Consequently, by either scaling the data or the bias, it may be possible to compensate for situations where a frame is not $\alpha$-rectifying on $K$ but is on $K^{\prime}\subsetneq K$. The following examples demonstrate that compensation can also be reached by other restrictions.
Figure 2: Left: The frame composed of the standard basis and its negative elements is 00-rectifying on $\mathbb{R}^{2}$. Mid: The standard basis is 00-rectifying on +$\mathbb{R}^{2}_{+}$ and $(-1)$-rectifying on $\mathbb{B}$. Right: The triangle frame is $(-\frac{1}{2})$-rectifying on $\mathbb{B}$, but never on $\mathbb{R}^{2}$ since there will always be cones where only one element is active (lighter areas).
Example 1.
A basis can never be $\alpha$-rectifying on $\mathbb{R}^{n}$ for any $\alpha$. However, the standard basis for $\mathbb{R}^{n}$ is $\alpha^{\prime}$-rectifying on +$\mathbb{R}^{n}_{+}$ for $\alpha^{\prime}\equiv 0$.
Example 2.
The frame
$\Phi_{3}=\left(\begin{pmatrix}0\\ 1\end{pmatrix},\begin{pmatrix}-\nicefrac{{\sqrt{3}}}{{2}}\\ -\nicefrac{{1}}{{2}}\end{pmatrix},\begin{pmatrix}\nicefrac{{\sqrt{3}}}{{2}}\\ -\nicefrac{{1}}{{2}}\end{pmatrix}\right)$
is not $\alpha$-rectifying on $\mathbb{R}^{n}$ for any $\alpha$ since $m=3<4=2n$. However, by a geometric argument (Figure 2), it is easy to see that $\Phi$ is $\alpha$-rectifying on $\mathbb{B}=\mathbb{B}_{1}$ for $\alpha\equiv-\frac{1}{2}$.
This makes clear that it is essential to select the domain carefully if we want to effectively study the injectivity behavior of the associated ReLU layer. This leads us to Question Q1.
3.1 Maximal Domain
We aim to identify the maximal domain $K$ for a frame $\Phi$ and a bias $\alpha$. This provides a comprehensive characterization of the $\alpha$-rectifying property, elucidating the injectivity of $C_{\alpha}$ from a geometric point of view.
Recall that for $i\in I$ and $\alpha\in\mathbb{R}^{m}$ we denote the closed affine half-space where the frame element $\phi_{i}$ is active for $\alpha$ by
$\Omega_{i}^{\alpha}=\{x\in\mathbb{R}^{n}:\langle x,\phi_{i}\rangle\geq\alpha_{i}\}.$ (24)
Extending the intuition from the example in Figure 1, we find that any frame will be $\alpha$-rectifying on the intersection of sufficiently many $\Omega_{i}^{\alpha}$’s. When restricting to minimal spanning sets, i.e., basis, we find that this is characterizing.
Theorem 4 (Maximal domain).
Let $\Phi\subset\mathbb{R}^{n}$ be a frame and $\alpha\in\mathbb{R}^{m}$. The maximal domain where $\Phi$ is $\alpha$-rectifying is given by
$\mathcal{K}_{\alpha}^{*}=\bigcup_{\begin{subarray}{c}J\subseteq I\\ \Phi_{J}\ \text{basis}\end{subarray}}\bigcap_{i\in J}\Omega_{i}^{\alpha}.$ (25)
In other words, $\Phi$ is $\alpha$-rectifying on $K$ if and only if $K\subseteq\mathcal{K}_{\alpha}^{*}$.
Proof.
Let $x\in K$. Assuming $\Phi$ to be $\alpha$-rectifying on $K$, then $\Phi_{I_{x}^{\alpha}}$ is a frame. Since every frame contains a basis, there is $L\subseteq I_{x}^{\alpha}$ such that the sub-collection $\Phi_{L}$ is a basis. Clearly, $\langle x,\phi_{i}\rangle\geq\alpha_{i}$ still holds for all $i\in L$, hence, $x\in\mathcal{K}_{\alpha}^{*}$.
For the converse direction, by definition of $\mathcal{K}_{\alpha}^{*}$ there is $M$ with $\langle x,\phi_{i}\rangle\geq\alpha_{i},\;x\in\mathcal{K}_{\alpha}^{*}$ for all $i\in M$ such that $\Phi_{M}$ is a basis. Since $M\subseteq I_{x}^{\alpha}$ must hold, $\Phi_{I_{x}^{\alpha}}$ is a frame. ∎
Figure 3: The dark areas in the left and mid picture indicate the maximal domains $\mathcal{K}_{\alpha}^{*}$ for the triangle frame and $\alpha\equiv 0$ (left), and a random frame on $\mathbb{S}$ with random bias (right). The right illustration corresponds to Example 4, where we point out how $K_{2}=\{x\in K:2\in J^{*}(x)\}$ looks like, where $J^{*}(x)$ is the most correlated basis for $x$.
When applying Theorem 4 to the closed ball $\mathbb{B}_{r}$ we find that a normalized frame $\Phi\subset\mathbb{S}$ is $\alpha$-rectifying on $\mathbb{B}_{r}$ if and only if
$r\leq\inf_{x\in\mathbb{R}^{n}\setminus\mathcal{K}_{\alpha}^{*}}\|x\|.$
Another consequence of the theorem is that for $\alpha\equiv 0$ we have that $\Phi$ is 00-rectifying on $\mathbb{B}_{r}$, $r>0$ if and only if $\Phi$ is 00-rectifying on $\mathbb{R}^{n}$. Hence, in this setting, checking a small neighborhood around the origin is already sufficient for the entire space. As an illustration that this does not hold for $\alpha\not\equiv 0$, we refer to the right plot in Figure 3 for an example.
The characterization in Theorem 4 further allows extending the implication from the $\alpha$-rectifying property to the injectivity of $C_{\alpha}$ (Theorem 2) to domains that are not open or convex, such as the sphere $\mathbb{S}$, the donut $\mathbb{D}_{r,s}$, and discrete data sets.
Corollary 2.
Let $\Phi=(\phi_{i})_{i\in I}\subset\mathbb{R}^{n}$, $\alpha\in\mathbb{R}^{m}$ and $\emptyset\neq K\subseteq O\subseteq\mathcal{K}_{\alpha}^{*}$ for $O$ open or convex. If $\Phi$ is $\alpha$-rectifying on $K$, then $C_{\alpha}$ is injective on $K$.
Proof.
Since $\Phi$ is $\alpha$-rectifying on $O\subseteq\mathcal{K}_{\alpha}^{*}$ open or convex, $C_{\alpha}$ is injective on $O$, hence also on $K\subseteq O$. ∎
It should be noted that, in general, the set $\mathcal{K}_{\alpha}^{*}$ is not open or convex. We can use this to demonstrate that a violation of the assumptions on $K$ in Theorem 2 (i.e., not open and not convex) indeed leads to the conclusion that $C_{\alpha}$ is not injective.
Example 3.
Consider the frame $\Phi_{3}$ defined in Example 2, and let
$x_{1}=\begin{pmatrix}\nicefrac{{1}}{{2}}\\ \nicefrac{{\sqrt{3}}}{{2}}\end{pmatrix},\;x_{2}=\begin{pmatrix}-\nicefrac{{1}}{{2}}\\ \nicefrac{{\sqrt{3}}}{{2}}\end{pmatrix}.$
See Figure 3 (left). For $\alpha\equiv 0$ we have that $x_{1},x_{2}\in\mathcal{K}_{\alpha}^{*}$ but also that $C_{\alpha}x_{1}=C_{\alpha}x_{3}$. Hence, by Theorem 4, $\Phi$ is $\alpha$-rectifying on $\mathcal{K}_{\alpha}^{*}$ but $C_{\alpha}$ is not injective on $\mathcal{K}_{\alpha}^{*}$.
An analog example can be constructed for the standard basis, $\alpha\equiv 0$, and $K=\mathbb{B}_{r}^{+}$ (Figure 2 mid). The characterization via the maximal domain provides a geometric intuition that reveals a natural trade-off, where the bias and domain affect each other inversely as
$\alpha^{\prime}\geq\alpha\quad\Leftrightarrow\quad\mathcal{K}_{\alpha^{\prime}}^{*}\subseteq\mathcal{K}_{\alpha}^{*}.$
This trade-off is not sharp for domains that are not maximal. Yet, it shows us that maximizing $\alpha$ can provide us with the same information from another perspective. With this, we proceed to answer Question Q2.
3.2 Maximal Bias
We aim to construct a maximal bias for a given frame $\Phi$ and domain $K$. This provides a characterization of the $\alpha$-rectifying property which is particularly interesting for verifying it in applications as it is straightforward to implement numerically.
Our approach to this is to decompose a frame $\Phi$ into sub-frames with highly correlated frame elements and identify the smallest analysis coefficients among all points $x\in K$ associated with these sub-frames. We present two approaches for such a decomposition. Approach A is based on the most correlated elements of $\Phi$ to each $x\in K$, and Approach B is based on the vertex-facet configuration of the inscribing polytope associated with $\Phi$. Corresponding algorithmic solutions are provided along with the theoretical results.
Approach A: Most Correlated Bases
As a preliminary stage to constructing a maximal bias, we construct a maximal constant bias via the least correlated frame element in the most correlated basis among all $x\in K$. This construction is similar to the one for the critical saturation level in [2]. We denote by $\overrightarrow{r}\in\mathbb{R}^{m}$ the constant vector with entries $r\in\mathbb{R}$.
Proposition 2.
Let $\Phi\subset\mathbb{R}^{n}$ be a frame and $K\subseteq\mathbb{R}^{n}$. The maximal constant bias for $\Phi$ and $K$ is given by $\overrightarrow{\alpha_{c}}$ with
$\alpha_{c}=\inf_{x\in K}\underset{\begin{subarray}{c}J\subseteq I\\ \Phi_{J}\ \text{basis}\end{subarray}}{\max}\ \min_{j\in J}\ \langle x,\phi_{j}\rangle.$ (26)
In other words, $\Phi$ is $\overrightarrow{r}$-rectifying on $K$ if and only if $r\leq\alpha_{c}$.
Proof.
First, we show that $\Phi$ is $\overrightarrow{\alpha_{c}}$-rectifying. Let $x\in K$ then there is a basis $\Phi_{J(x)}$ such that for all $j\in J(x)$
$\langle x,\phi_{j}\rangle\geq\min_{j\in J(x)}\langle x,\phi_{j}\rangle\geq\alpha_{c}.$
Since $\Phi_{J(x)}$ is a frame, $\Phi$ is $\overrightarrow{\alpha_{c}}$-rectifying.
Now let $r\in\mathbb{R}$ and assume that $\Phi$ is $\overrightarrow{r}$-rectifying. For any $x\in K$ we deduce
$r\leq\inf_{x\in K}\min_{j\in I_{x}^{\overrightarrow{r}}}\langle x,\phi_{j}\rangle\leq\inf_{x\in K}\underset{\begin{subarray}{c}j\in J\subseteq I\\ \Phi_{J}\ \text{basis}\end{subarray}}{\min}\langle x,\phi_{j}\rangle\leq\inf_{x\in K}\underset{\begin{subarray}{c}J\subseteq I\\ \Phi_{J}\ \text{basis}\end{subarray}}{\max}\ \min_{j\in J}\ \langle x,\phi_{j}\rangle=\alpha_{c}.$ (27)
To construct a (non-constant) maximal bias vector we introduce the notion of a most correlated basis for $x\in K$. We define this formally as follows.
Definition 3.
Let $\Phi$ be a frame and $x\in K$. If $J^{*}(x)\subseteq I$ satisfies that for all $J\subseteq I$ such that $\Phi_{J}$ is a basis it holds that
$\min_{j\in J}\ \langle x,\phi_{j}\rangle\leq\min_{j\in J^{*}(x)}\ \langle x,\phi_{j}\rangle$ (28)
we call $\Phi_{J^{*}(x)}$ a most correlated basis for $x$. We say that $\Phi$ has a unique most correlated basis everywhere if $J^{*}(x)$ is unique for every $x\in K$.
Alternatively, we can say that $J^{*}(x)$ maximizes the functional
$\alpha(x)=\underset{\begin{subarray}{c}J\subseteq I\\ \Phi_{J}\ \text{basis}\end{subarray}}{\max}\ \min_{j\in J}\ \langle x,\phi_{j}\rangle.$ (29)
We will restrict ourselves to frames where the most correlated basis is unique everywhere. Similar to the full-spark assumption, in a numerical setting this is a mild condition since for any frame there is an arbitrarily small perturbation such that the resulting perturbed frame fulfills the requirement (c.f. Lemma 2). If the frame is full-spark then the most correlated basis for $x$ is given by the collection of the $n$ frame elements which have the largest frame coefficients with $x$. A random frame fulfills this condition with probability one. Under this assumption, we can partition $K$ uniquely into subsets $K_{i}$ each associated with a frame element that belongs to a most correlated basis, i.e.,
$K_{i}=\{x\in K:i\in J^{*}(x)\}$ (30)
for every $i\in I$. Since every $x\in K$ has a most correlated basis $\bigcup_{i\in I}K_{i}=K$ indeed holds. The right picture in Figure 3 illustrates the set $K_{i}$ in $\mathbb{R}^{2}$, and the right plot in Figure 4 illustrates the partition of $\mathbb{S}$ in $\mathbb{R}^{3}$ into $K_{i}$’s. By minimizing the frame coefficients of $x\in K_{i}$ and the $\phi_{i}$’s for any $i\in I$ similar to (26) we obtain a large bias such that $\Phi$ possesses the rectifying property with that bias - but is in general not maximal.
Proposition 3.
Let $\Phi\subset\mathbb{R}^{n}$ be a frame and $K\subseteq\mathbb{R}^{n}$. If $\alpha^{\flat}_{K}$ is given as
$\left(\alpha^{\flat}_{K}\right)_{i}=\inf_{x\in K_{i}}\langle x,\phi_{i}\rangle,$ (31)
then $\Phi$ is $\alpha^{\flat}_{K}$-rectifying on $K$.
Proof.
Let $x\in K$ then $\langle x,\phi_{j}\rangle\geq(\alpha^{\flat}_{K})_{j}$ for all $j\in J^{*}(x)$. Since $\Phi_{J^{*}(x)}$ is a basis, $\Phi$ is $\alpha^{\flat}_{K}$-rectifying on $K$. ∎
Note that if a $K_{i}$ is empty then $\phi_{i}$ is never contained in a most correlated basis, hence, $\phi_{i}$ is irrelevant for the rectifying property of $\Phi$.
Although $\alpha^{\flat}_{K}$ gives a simple and intuitive indication of the critical bias level, there is still room for increasing this bias while maintaining the rectifying property. We can do this by minimizing over the sets
$K_{i}^{\sharp}=K_{i}\setminus\Bigg{(}\bigcap_{y\in K_{i}}\bigcup_{j\notin J^{*}(y)}\Omega_{j}^{\alpha^{\flat}_{K}}\Bigg{)}.$ (32)
instead of $K_{i}$. The sets $K_{i}^{\sharp}$ contain all points $x$ such that $i\in J^{*}(x)$ and no element outside the most correlated basis is active for $x$ and $\alpha^{\flat}_{K}$. Since $K_{i}^{\sharp}\subseteq K_{i}$ the bias that we get from minimizing over the $K_{i}^{\sharp}$ will be larger than $\alpha^{\flat}_{K}$. Note that if $K_{i}\subseteq\big{(}\bigcap_{y\in K_{i}}\bigcup_{j\notin J^{*}(y)}\Omega_{j}^{\alpha^{\sharp}_{K}}\big{)}$ then whenever $\phi_{i}$ belongs to the most correlated basis for an $x$, there is also another active element from outside the most correlated basis. If this happens, the frame element $\phi_{i}$ can be interpreted as being redundant in the sense that it can be removed from $\Phi$ while preserving the rectifying property. We give an example of such a pathetic situation.
Example 4.
For the frame
$\Phi=\left(\begin{pmatrix}1\\ 0\end{pmatrix},\begin{pmatrix}\nicefrac{{1}}{{\sqrt{2}}}\\ \nicefrac{{1}}{{\sqrt{2}}}\end{pmatrix},\begin{pmatrix}0\\ 1\end{pmatrix},\begin{pmatrix}-1\\ 0\end{pmatrix},\begin{pmatrix}0\\ -1\end{pmatrix}\right)$
and $K=\mathbb{B}$ we have that $\alpha_{\mathbb{B}}^{\flat}\equiv 0$ and
$K_{2}=\left\{s\cdot\begin{pmatrix}\cos{t}\\ \sin{t}\end{pmatrix},t\in\left[-\frac{3\pi}{8},\frac{3\pi}{8}\right],s\in\left[0,1\right]\right\}.$
The right picture in Figure 3 shows this setting. For every $x\in K_{2}$ there is an element outside the most correlated basis that is additionally active. It follows that $K_{2}^{\sharp}=\emptyset$. Hence, $\phi_{2}$ is irrelevant for the rectifying property of $\Phi$.
To facilitate the formulation of the maximality statement we will make the reasonable and mild assumption that $K_{i}^{\sharp}\neq\emptyset$ for all $i\in I$. Otherwise, a maximal bias will not exist in general. The following theorem presents the main statement on the maximal bias, and is the counterpart to Theorem 4.
Theorem 5 (Maximal bias).
Let $\Phi\subset\mathbb{R}^{n}$ be a frame that has a unique most correlated basis everywhere, $K\subseteq\mathbb{R}^{n}$, and $K_{i}^{\sharp}\neq\emptyset$ for all $i\in I$. If $\alpha^{\sharp}_{K}$ is given as
$\left(\alpha^{\sharp}_{K}\right)_{i}=\inf_{x\in K_{i}^{\sharp}}\langle x,\phi_{i}\rangle,$ (33)
then $\Phi$ is $\alpha$-rectifying on $K$ if and only if $\alpha\leq\alpha^{\sharp}_{K}$.
Proof.
For the implication direction, we have that $\langle x,\phi_{j}\rangle\geq(\alpha^{\sharp}_{K})_{j}$ for all $j\in J^{*}(x)$. Since all $K_{j}^{\sharp}$ are non-empty, $\Phi$ is $\alpha^{\sharp}_{K}$-rectifying on $K$.
We show the converse direction by counterposition. Let $i\in I$ and assume that $\Phi$ is $\alpha$-rectifying on $K$ where $\alpha$ is given by $\alpha_{i}=(\alpha^{\sharp}_{K})_{i}+\varepsilon$ for some $\varepsilon>0$ and $\alpha_{j}=(\alpha^{\sharp}_{K})_{j}$ for $j\neq i$. By definition of the infimum in (33), the set $K_{i}\sharp$ (32), and the construction and uniqueness of the most correlated basis (29) there is $x_{0}\in K_{i}$ such that
$\langle x_{0},\phi_{i}\rangle<\left(\alpha^{\sharp}_{K}\right)_{i}+\varepsilon\qquad\text{and}\qquad\langle x_{0},\phi_{j}\rangle<\left(\alpha^{\sharp}_{K}\right)_{j}$ (34)
for all $j\notin J^{*}(x_{0})$. This implies that the active coordinates for $x_{0}$ and $\alpha$ are exactly given by
$I_{x_{0}}^{\alpha}=J^{*}(x_{0})\setminus\{i\}.$
As a consequence, $\Phi_{I_{x_{0}}^{\alpha}}$ does not form a frame. Since $x_{0}$ was the minimizer of $\left(\alpha^{\sharp}_{K}\right)_{i}$ we can deduce that (34) holds for all $x\in K_{i}$. Therefore, $\Phi$ is not $\alpha$-rectifying on $K_{i}$. Since $K=\bigcup_{i\in I}=K$, we obtain the desired counterposition argument. ∎
Summarizing, Proposition 3 provides a simple and intuitive sufficient condition for a frame $\Phi$ to possess the rectifying property on $K$ via $\alpha^{\flat}_{K}$, whereas Theorem 5 provides a full characterization via the more complicated $\alpha^{\sharp}_{K}$ under some additional mild conditions. For an algorithmic approach, we focus on $\alpha^{\flat}_{K}$.
Algorithmic approach. We show how $\alpha_{K}^{\flat}$ can be approximated numerically via sampling. Let $X_{N}=(x_{k})_{k=1}^{N}\subset K$ be a sequence of $N$ samples in $K$. By iteratively updating the values of the bias estimation corresponding to the most correlated basis of the current sample, in analogy to (31), we obtain an approximation of $\alpha_{K}^{\flat}$ through $X_{N}$. To measure the approximation error quantitatively, we use the Euclidean covering radius (or ”mesh norm“) of $X_{N}$ for $K$ [13], defined by
$\rho(X_{N};K)=\sup_{x\in K}\min_{1\leq i\leq N}\|x-x_{i}\|.$ (35)
Theorem 6 (Sampling-based Bias Estimation).
Let $X_{N}=\{x_{i}\}_{i=1}^{N}\subset K$ and $\Phi\subset\mathbb{S}$ be a normalized full-spark frame. Choose $\alpha^{(0)}\in\mathbb{R}^{m}$ and let for any $1\leq k\leq N$
$(\alpha^{(k)})_{i}=\min\left\{\langle x_{k},\phi_{i}\rangle,\alpha^{(k-1)}_{i}\right\},$ (36)
for all $i\in J^{*}(x_{k})$. Then $\Phi$ is $\alpha^{(N)}$-rectifying on $X_{N}$. Moreover, $\Phi$ is $(\alpha^{(N)}-\rho(X_{N};K))$-rectifying on $K$.
Note that if the elements in $\Phi$ are not normalized, we can extend the statement by including the norms $w=(\|\phi_{i}\|)_{i\in I}$ such that $\Phi$ is $(\alpha^{(N)}-\rho(X_{N};K)\cdot w)$-rectifying on $K$. If the frame elements of $\Phi$ lie in $K$, a good initialization is starting the algorithm with those directly as samples. Otherwise, we may set $\alpha^{(0)}\equiv\infty$.
Theorem 6.
Let $x\in K$, then there is $x_{i}\in X_{N}$ with $\|x-x_{i}\|\leq\rho(X_{N};K)$. Furthermore, for every $j\in I_{x_{i}}^{\alpha^{(N)}}$ it holds that
$\begin{split}(\alpha^{(N)})_{j}&\leq\langle x_{i},\phi_{j}\rangle\\ &=\langle x_{i}-x,\phi_{j}\rangle+\langle x,\phi_{j}\rangle\\ &\leq\|x-x_{i}\|+\langle x,\phi_{j}\rangle\\ &\leq\rho(X_{N};K)+\langle x,\phi_{j}\rangle.\end{split}$ (37)
By rearranging (37) it follows that $\langle x,\phi_{j}\rangle\geq(\alpha^{(N)})_{j}-\rho(X_{N};K)$. We deduce that $I_{x_{i}}^{\alpha^{(N)}}\subseteq I_{x}^{\alpha^{(N)}-\rho(X_{N};K)}$ for all $x\in K$. Clearly, $\Phi$ is $\alpha^{(N)}$-rectifying on $X_{N}$ by construction. Using (37), the second claim follows immediately. ∎
Implementing the algorithm described in Theorem 6 in practice can be done via a Monte-Carlo approach, where $X_{N}$ is a collection of $N$ random samples on $K$ w.r.t. some probability measure on $K$. We will demonstrate numerical experiments in Section 3.3. Intuitively, we want to have a measure that guarantees a small covering radius $\rho(X_{N};K)$ for large $N$, i.e., that $\alpha^{(N)}$ converges to $\alpha_{K}^{\flat}$ in probability. The uniform distribution on $\mathbb{B}_{r}$ is a possible example. If $X_{N}$ are i.i.d. uniform samples on $\mathbb{S}$ the expectation of $\rho(X_{N};\mathbb{S})$ is given by
$\mathbb{E}[\rho(X_{N};\mathbb{S})]\asymp\left(\frac{\log(N)}{N}\right)^{\frac{1}{n}}.$ (38)
Explicit tail-bound estimates for the probability distribution of $\rho(X_{N};\mathbb{S})$ are derived in [27]. Unfortunately, this indicates that in high dimensions it becomes infeasible to handle the covering radius only by increasing the number of test samples $N$. Hence, it appears necessary to construct the sampling sequence $X_{N}$ in a more structured way, e.g., by quasi Monte-Carlo methods [6], or by sampling directly from the distribution of the dataset. With the latter approach, the number of sampling points for a good approximation of $\alpha_{K}^{\flat}$ can potentially be reduced, and the injectivity of the ReLU-layer is ensured on a domain that is tailored to the dataset.
Approach B: Facets of the Inscribing Polytope
We obtain another natural decomposition of $\Phi$ into sub-frames with high correlation via the convex polytope that arises as the set of all convex linear combinations of the elements in $\Phi$, also called the inscribing polytope of $\Phi$ [28], given by
$P_{\Phi}=\{x\in\mathbb{R}^{n}:x=\sum_{i\in I}c_{i}\cdot\phi_{i},c_{i}\geq 0,\sum_{i\in I}c_{i}=1\}.$ (39)
Any non-empty intersection of $P_{\Phi}$ with an affine half-space such that none of the interior points of $P_{\Phi}$ (w.r.t. the induced topology on $P_{\Phi}$) lie on its boundary is called a face of $P_{\Phi}$ [32]. The 00-dimensional faces of $P_{\Phi}$ are known as vertices and the $(n-1)$-dimensional faces are called facets. Motivated by the discussion on normalized frames in Section 2.2, we assume $\Phi\subset\mathbb{S}$ here. This implies that the set of vertices of $P_{\Phi}$ always coincides with the frame elements of $\Phi$. (This is generally the case if the elements in $\Phi$ lie in a strictly convex set.) Then, every facet is a convex polytope where the vertices coincide with a sub-collection of $\Phi$, and every element occurs as vertex at least once. For a facet $F$, we denote the index set corresponding to its vertices by
$I_{F}=\{i\in I:\phi_{i}\in F\}.$ (40)
See Figure 5. The following property is key, stated and proven in [16].
Lemma 5.
Let $F$ be a facet of $P_{\Phi}$. If $0\notin F$, then $\Phi_{I_{F}}$ is a frame.
Hence, the facets of $P_{\Phi}$ yields a natural decomposition into sub-frames $\Phi_{I_{F}}$ of $\Phi$. Moreover, for any facet $F$, there is $a\in\mathbb{R}^{n}$, $a\neq 0$ and $b\in\mathbb{R}$ such that $F=\{x\in P_{\Phi}:\langle a,x\rangle=b\}$, and therefore,
$\langle a,\phi_{k}\rangle$ $=b,\;\text{for }k\in I_{F},$
$\langle a,\phi_{\ell}\rangle$ $<b,\;\text{for }\ell\notin I_{F}.$
This indicates that the vertices of any facet form a frame which is highly correlated to all points “close” to the facet. In particular, $\Phi_{I_{F}}$ is the most correlated basis for the normal vector $a$.
Figure 4: The three subplots show different decompositions of the sphere in $\mathbb{R}^{3}$. The black dots indicate the $m=12$ frame elements on $\mathbb{S}$. From left to right: The facets $F_{j}$ of the inscribing polytope $P_{\Phi}$, the associated spherical caps $F_{j}^{\mathbb{S}}=\operatorname{cone}(F_{j})\cap\mathbb{S}$, and the spherical patches associated to different most correlated bases obtained by $J^{*}(x)$. While the facets provide a very intuitive and simple grouping into sub-frames, the grouping via the most correlated bases minimizes the correlation.
Now, analog to the decomposition of $K$ using most correlated bases (30) we can partition $K$ into facet-specific subsets $F_{j}^{K}\subseteq K$ for every facet $F_{j}$ of $P_{\Phi}$ (according to some enumeration of the facets)
$K=\bigcup_{j}F_{j}^{K}.$ (41)
A natural and practical approach in this setting is a conical partition, i.e.,
$F_{j}^{K}=\operatorname{cone}(F_{j})\cap K=\{x\in K:x=cy\;,y\in F_{j},c\geq 0\}.$ (42)
Figure 4 shows the facets $F_{j}$ of the inscribing polytope (left), and the decomposition of $\mathbb{S}$ into spherical caps $F_{j}^{\mathbb{S}}$ (mid) for a frame $\Phi$ with random i.i.d. elements on $\mathbb{S}$.
To guarantee that (41) holds and that 00 is not contained in any of the facets (see Lemma 5) we have to assume that 00 lies in the interior of $P_{\Phi}$ (w.r.t. the topology in $\mathbb{R}^{n}$). This property was introduced in [4, Definition 1], where in this case $\Phi$ is called omnidirectional. Equivalently, we can say that there is no half-space containing $\Phi$, which can be easily verified numerically via convex optimization, see appendix [4]. This assumption is natural if $K$ is centered around the origin. Moreover, every non-omnidirectional frame can be made omnidirectional by including the vector that arises as the mean of all frame elements, see appendix.
Remark 1.
Depending on the properties of $K$ (e.g., not centered around the origin), alternative constructions of $F_{j}^{K}$ might be more appropriate than the one in (42). Also, the assumption of omnidirectionality might become obsolete and needs to be adapted accordingly. In this work, we restrict ourselves to the described setting.
Making use of Lemma 5 and the partitioning of $K$ in (41), we identify the minimal analysis coefficient $\langle y,\phi_{i}\rangle$ that can occur for $y$ on each $F_{j}^{K}$ that contains the vertex $\phi_{i}$. This guarantees that the sub-frame $\Phi_{I_{F_{j}}}$ is active for any $x\in F_{j}^{K}$. We call the procedure polytope bias estimation (PBE), as in [16]. The following theorem generalizes the results in [16] from $\mathbb{B}_{r}$ and $\mathbb{B}_{r}^{+}$ to general bounded $K$.
Theorem 7 (Polytope Bias Estimation).
Let $\Phi\subset\mathbb{S}$ be an omnidirectional
frame and $K$ bounded. Then $\Phi$ is $\alpha_{K}^{\Delta}$-rectifying on $K\subseteq\mathbb{R}^{n}$ with $\alpha_{K}^{\Delta}$ given by
$\left(\alpha_{K}^{\Delta}\right)_{i}=\inf_{\begin{subarray}{c}y\in F_{j}^{K}\\ j:\phi_{i}\in F_{j}\end{subarray}}\langle y,\phi_{i}\rangle.$ (43)
Proof.
Since $\Phi$ is omnidirectional, for any $x\in K$ there is a facet $F_{j}$ such that $x\in F_{j}^{K}$. It follows from (43) that $\langle x,\phi_{i}\rangle\geq\left(\alpha_{K}^{\Delta}\right)_{i}$ for all $i\in I_{F_{j}}$. By Lemma 5, $\Phi_{I_{F_{j}}}$ is a frame, hence, $\Phi$ is $\alpha_{K}^{\Delta}$-rectifying on $K$. ∎
This procedure naturally respects the geometry of the frame and provides an intuitive way of estimating a large bias vector such that the associated ReLU-layer is injective.
Algorithmic approach
The advantage of the PBE is that for standard domains $K$ it is easy to compute $\alpha_{K}^{\Delta}$ via simple linear programs. In the following proposition, we formulate the PBE for five prototypical domains.
Proposition 4.
Let $\Phi\subset\mathbb{S}$ be an omnidirectional frame. The following holds.
  1. (i)
    $\Phi$ is $\alpha_{\Phi}^{\Delta}$-rectifying on the boundary of the polytope $\partial P_{\Phi}=\bigcup_{j}F_{j}$ with $\alpha_{\Phi}^{\Delta}$ given by
    $\left(\alpha_{\Phi}^{\Delta}\right)_{i}=\min_{\begin{subarray}{c}\ell\in I_{F_{j}}\\ j:\phi_{i}\in F_{j}\end{subarray}}\langle\phi_{\ell},\phi_{i}\rangle.$
  2. (ii)
    $\Phi$ is $\alpha_{\mathbb{S}}^{\Delta}$-rectifying on the sphere $\mathbb{S}$ with $\alpha_{\mathbb{S}}^{\Delta}$ given by
    $\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i}=\min\{\min_{\begin{subarray}{c}y\in F_{j}^{\mathbb{S}}\\ j:\phi_{i}\in F_{j}\end{subarray}}\langle y,\phi_{i}\rangle,\left(\alpha_{\Phi}^{\Delta}\right)_{i}\}.$
    The inner minima are the solutions of convex linear programs. In particular, they are equal to $\left(\alpha_{\Phi}^{\Delta}\right)_{i}$ whenever they are non-negative.
  3. (iii)
    $\Phi$ is $(r^{-1}\cdot\alpha_{\mathbb{B}}^{\Delta})$-rectifying on the donut $\mathbb{D}_{r,s}$ for $0\leq s<r$ with $\alpha_{\mathbb{B}}^{\Delta}$ given by
    $\left(\alpha_{\mathbb{B}}^{\Delta}\right)_{i}=\min\{s,\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i}\}.$
    The case $s=0$ yields a bias estimation for the closed ball $\mathbb{B}_{r}$ (Figure 5).
  4. (iv)
    Let $J^{+}=\{j\in I:F_{j}\cap\mathbb{R}^{n}_{+}\neq\emptyset\}$ and $I^{+}=\bigcup_{j\in J^{+}}I_{F_{j}}$ then $\Phi$ is $(r^{-1}\cdot\alpha_{\mathbb{B}^{+}}^{\Delta})$-rectifying on the non-negative part of the ball ${\mathbb{B}^{+}_{r}}$ with $\alpha_{\mathbb{B}^{+}}^{\Delta}$ given by
    $\left(\alpha_{\mathbb{B}^{+}}^{\Delta}\right)_{i}=\begin{cases}\left(\alpha_{\mathbb{B}}^{\Delta}\right)_{i}&\text{ for }i\in I^{+}\\ s_{i}&\text{ else,}\end{cases}$ (44)
    where $s_{i}\in\mathbb{R}$ is arbitrary.
  5. (v)
    If $\alpha_{\Phi}^{\Delta}\geq 0$, then $\left(s\cdot\phi_{i}\right)_{i\in I}$ for $s\geq 0$ is $\left(s\cdot\alpha_{\Phi}^{\Delta}\right)$-rectifying on $(\mathring{\mathbb{B}}_{s})^{\mathrm{C}}=\mathbb{R}^{n}\setminus\mathring{\mathbb{B}}_{s}$.
Proof.
The points $(i)-(iv)$ are direct consequences of [16, Theorem 4.4 and Theorem 4.6], where detailed proofs can be found. To show $(v)$, note that $\left(s\cdot\phi_{i}\right)_{i\in I}$ has the same combinatorial facet structure as $\Phi$, hence, is still omnidirectional. Therefore,
$(\mathring{\mathbb{B}}_{s})^{C}=\{y\in\mathbb{R}^{n}:y=st\cdot x,x\in\mathbb{S},t\geq 1\}.$
The statement follows by $\langle st\cdot x,\phi_{i}\rangle\geq s\cdot\left(\alpha_{\Phi}^{\Delta}\right)_{i}$ for $t\geq 1$ and $s\geq 0$. ∎
Figure 5: For the Icosahedron frame we have that $x\in F_{j}\Leftrightarrow I_{F_{j}}=J^{*}(x)$ (left). For less regular frames, this does not hold anymore (mid). The right picture illustrates the PBE on $\mathbb{B}$ for the Icosahedron frame. To get $\left(\alpha_{\mathbb{B}}^{\Delta}\right)_{2}$ the infima are taken over the points in the conical parts (dark gray area) for all adjacent facets of $\phi_{2}$ (blue).
For all settings described above we provide the pseudo-code for implementing the corresponding optimization in the appendix. Other than the sampling-based approach, this bias estimation procedure yields a deterministic sufficient condition for the $\alpha$-rectifying property, which is also necessary only in special situations. In general, there are two properties of the inscribing polytope $P_{\Phi}$ that affect the PBE in a way that it does not yield a maximal bias.
  • (1)
    The more vertices a facet has, the smaller the infima in (43) become, i.e., the estimation is worse. In the case where all facets have exactly $n$ vertices, $P_{\Phi}$ is called simplicial. Fortunately, this is the generic case. For example, it is known that a polytope with vertices that are i.i.d. uniform samples on $\mathbb{S}$ is simplicial with probability one [8].
  • (2)
    The decomposition of $K$ using cones in (41) is natural and convenient for implementation, but leads to a sub-optimal partition if $P_{\Phi}$ is geometrically very irregular, i.e., the sizes of the facets differ a lot. In such a scenario, for $x\in F_{j}^{K}$ where $F_{j}$ is a very large facet, there might be a smaller neighboring facet $F_{k}$ which provides larger analysis coefficients and, hence, a better estimation. For an illustration of such a situation see the center plot in Figure 5.
For frames with inscribing polytopes that are simplicial and regular, we can show that the PBE indeed yields a maximal bias. We set $K=\mathbb{S}$.
Lemma 6.
Let $\Phi\subset\mathbb{S}$ such that $P_{\Phi}$ is simplicial and for all $i\in I$ it holds that $\langle\phi_{i},\phi_{k}\rangle=\langle\phi_{i},\phi_{\ell}\rangle$ for all $\phi_{k},\phi_{\ell}$ sharing a facet with $\phi_{i}$. Then $\Phi$ is $\alpha$-rectifying on $\mathbb{S}$ if and only if $\alpha\leq\alpha_{\mathbb{S}}^{\Delta}$.
Proof.
The assumption on $P_{\Phi}$ implies that $\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i}=\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{k}$ for all $i,k\in I$. Let $\Phi$ be $\alpha$-rectifying for $\alpha\in\mathbb{R}^{m}$ with $\alpha_{i^{*}}>\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i^{*}}$ for $i^{*}\in I$. There is a facet $F_{j}$ and $x^{*}\in F_{j}^{\mathbb{S}}$ with $\langle x^{*},\phi_{i^{*}}\rangle=\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i^{*}}$, implying that $\langle x^{*},\phi_{i^{*}}\rangle<\alpha_{i^{*}}$. Since $P_{\Phi}$ is simplicial $\Phi_{I_{F_{j}}\setminus\{{i^{*}}\}}$ is not a frame. Hence, $\alpha_{i}\leq\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i}$ must hold for all $i\in I$. ∎
In this sense, we expect the PBE to be very effective for frames that are evenly distributed on $\mathbb{S}$.
Remark 2.
The two bias estimation procedures described in Approach A and B are fundamentally linked. For each facet $F$ of $P_{\Phi}$ the hole radius of $F$ is defined as the Euclidean distance from the boundary to the center of its spherical cap $F^{\mathbb{S}}$ (42). The Euclidean covering radius (35) of $\Phi$ for $\mathbb{S}$ is the largest hole radius among all facets [30].
3.3 Numerical Experiments
For a given sampling sequence $X_{N}$ and a full-spark assumption of $\Phi$ the numerical implementation of the sampling-based bias estimation in Theorem 6 is straightforward. Similarly, there are convex hull algorithms available, such that the implementation of the PBE (in particular, the cases in Proposition 4) is straightforward, too. We provide pseudo-code for both approaches in the appendix. Concrete implementations in Python, together with the code for reproducing all experiments can be downloaded from the accompanying repository https://github.com/danedane-haider/Alpha-rectifying-frames.
Evolution Towards Injectivity (Approach A)
We demonstrate the basic functionality of the sampling-based algorithm in a random setting. We choose the frame $\Phi$ and the sampling sequences $X_{N}$ to consist of i.i.d. uniform samples on $\mathbb{B}$. For every step in the approximation we measure the proportion of samples from an unseen test sampling sequence $Y_{M}$ (also i.i.d. uniform on $\mathbb{B}$), for those $\Phi$ is $\alpha^{(k)}$-rectifying, i.e. the ReLU-layer is “point-wise” injective. Figure 6 shows the empirical mean and variance over $1000$ independent trials of this procedure for two different dimensions, $n=3$ (left) and $n=30$ (right), and three different redundancies ($2,3.3,$ and $9$), respectively. We observe that in some cases the associated ReLU-layer is injective already after a few hundred iterations. In other cases, it takes about $1000$ iterations. It is especially fast for ReLU-layers in low dimensions with low redundancy. We note that for draws of $\Phi$ which yield very unevenly distributed points on $\mathbb{B}$ (which happens more likely for layers in high dimensions with high redundancy) the iterative scheme struggles to update $\alpha^{(k)}$ efficiently, which results in the procedure taking particularly long. In general, all tested examples became fully injective reliably.
Figure 6: Per iteration $k$, the plots show the proportion of a test sample sequence $Y_{M}$ where the ReLU-layer with bias $\alpha^{(k)}$ is injective. Left: $n=3$. Right: $n=30$. Both for redundancies $2,3.3,$ and $9$. The ReLU-layers are becoming injective reliably after about $10^{4}$ iterations. The procedure is fastest for low dimensions and low redundancy.
Effect of Redundancy (Approach A)
We use approximations of $\alpha_{K}^{\flat}$ to verify the injectivity of ReLU-layers with random weights and biases systematically for different redundancies. We let both, the frame $\Phi$ and the sampling sequence $X_{N}$ contain i.i.d. standard normal points and compute $\alpha^{(N)}$ with $N=5\cdot 10^{5}$ for all redundancy settings where $2\leq n\leq 30$ and $n\leq m\leq 150$. By Theorem 6, we know that for every setting $\Phi$ is $\alpha^{(N)}$-rectifying on $X_{N}$. Inspired by the asymptotic expression for the covering radius in (38), we subtract a correcting term of $\rho^{*}(n,N)=0.05\cdot\left(\frac{\log(N)}{N}\right)^{\frac{1}{n}}$ to compensate for insufficient amount of sampling in higher dimensions. This yields that for every setting we have that $\Phi$ is $\left(\alpha^{(N)}-\rho^{*}(n,N)\right)$-rectifying on $\mathbb{B}$ with high probability. The factor $0.05$ was chosen experimentally.
To test if the ReLU-layer associated with one of the realizations of $\Phi$ is injective for a given bias $\alpha$ we have to verify that $\alpha\leq\alpha^{(N)}-\rho^{*}(n,N)$. We compare three settings for biases with i.i.d. normal values with mean zero and variances,
$(i)\ \sigma^{2}=0\qquad\quad(ii)\ \sigma^{2}=0.1\qquad\quad(iii)\ \sigma^{2}=1$
Figure 7 shows the results for the three settings from left to right. Setting $(i)$ is the same as the one where it was conjectured that the transition from non-injectivity to injectivity happens at a redundancy $q\in(6.6979,6.6981)$ [23]. Looking at Figure 7 (left) our approach supports this conjecture well. On the injectivity of random ReLU-layers with non-trivial bias (i.e., non-zero), there are no theoretical results in the literature so far, and our approach yields some insights here. For small variance (Setting $(ii)$) we observe that the clear boundary from the trivial bias blurs out. For the standard variance in setting $(iii)$ this behavior further intensifies. Note that changing the variances in $\Phi$ or $X_{N}$ has an equivalent effect. From these observations, we deduce that bias initialization in a ReLU-layer has a large impact on its injectivity.
Figure 7: The plots show the injectivity behavior of random ReLU-layers for different redundancies ($2\leq n\leq 30,n\leq m\leq 150$). The brightness encodes the proportion of the values of the given bias $\alpha$ that are smaller than the ones in $(\alpha^{(N)}-\rho^{*}(n,N))$ for $N=5\cdot 10^{5}$. For values of one (yellow), the ReLU-layer is injective on $\mathbb{B}$. In all settings, the samples in $X_{N}$ are i.i.d. standard normal and the bias i.i.d. normal with different variances. From left to right: $\sigma^{2}=0$, $\sigma^{2}=0.1$, and $\sigma^{2}=1$. For $\sigma^{2}=0$ we observe a very clear border from non-injective to injective at a redundancy of $6.7$ (solid magenta line) that aligns well with the theoretical results from the literature. For larger variance, the border blurs out quickly and prevents us from predicting clear statements about injectivity.
Approximation of a Maximal Bias (Approach A & B)
In the setting of Lemma 6 (i.e., $P_{\Phi}$ is simplicial and regular) we have shown that the PBE yields a maximal bias. This allows us to study the approximation of the sampling-based approach to the maximal bias $\alpha^{\mathbb{S}}$ by the PBE over the iterations. Figure 8 shows this in the example of the Tetrahedron frame, where the PBE yields $\alpha^{\mathbb{S}}\equiv-\frac{1}{\sqrt{3}}$. We plot $\|\alpha^{(k)}-\alpha^{\mathbb{S}}\|$ as $k$ increases and find that the approximation is very slow here, which emphasizes the superiority of the polytope approach in this setting.
Figure 8: Left: The inscribing polytope for the Tetrahedron frame. Right: The Euclidean distance of $\alpha^{(k)}$ to the maximal bias of the Tetrahedron frame $\alpha^{\mathbb{S}}$ on a log y-scale over $10^{5}$ iterations. In this example, where the inscribing polytope is a simplicial and regular polytope, the sampling-based bias estimation is sub-optimal, and the polytope approach already gives the maximal bias.
Remarks on Limitations and Area of Application
In general, both algorithms suffer from high dimensionality. For the sampling-based approach, the asymptotic behavior of the covering radius (38) indicates that it may become infeasible to reach a good approximation of a maximal bias only by increasing the number of samples. Yet, with some experimenting on the factor, we can use expression (38) to effectively compensate for insufficient sampling in high dimensions. The sampling-based approach has particularly high potential when injectivity is only required on specific data points of interest. In such a situation, the sampling set can be constructed in a data-driven way that respects the distribution of the data.
For the Polytope bias estimation, the numerical computation of the convex hull to obtain the vertex-facet incidences becomes computationally infeasible in high dimensions. A possible remedy is to use dimensionality reduction methods and do the bias estimation in a lower dimensional space. The benefits of the method are that injectivity follows deterministically and that it comes with a lot of intuition and can be studied further using more advanced tools from convex geometry. However, so far it is only applicable for the prototypical domains discussed in Proposition 4.
With this, we conclude the part of the paper that is concerned with the analysis of the injectivity of a ReLU-layer. The last chapter is dedicated to the actual reconstruction of the input from the output of an injective ReLU-layer.
4 Duality and Reconstruction
Assuming a ReLU-layer $C_{\alpha}$ to be injective, it is possible to construct an inverse mapping that can infer any input from the output. This extends to the possibility of synthesizing new data that correspond to arbitrary coefficient vectors in the image of the ReLU-layer, or studying the connection of single weights to the input by perturbing the corresponding output coefficient.
In [2], the authors propose to reconstruct the input from its saturated frame coefficients via a customly modified version of the frame algorithm [11]. Other than this approach, we propose to construct explicit perfect reconstruction formulas to obtain $x\in K$ from the output $C_{\alpha}x$ using the notion of duality from frame theory.
4.1 ReLU-synthesis
First, we recall the concept of a dual frame, which is closely tied to two operators. The first one is the synthesis operator associated with $\Phi=(\phi_{i})_{i\in I}$ which maps the frame coefficients back to the signal space, $\mathbb{R}^{n}$, as
$D:\mathbb{R}^{m}$ $\rightarrow\mathbb{R}^{n}$
$(c_{i})_{i\in I}$ $\mapsto\sum_{i\in I}c_{i}\cdot\phi_{i}.$
The realization of this operator is done via the multiplication by the transpose of the analysis matrix from the left, i.e., $D=C^{\top}$. The second one is the operator that arises as the concatenation of analysis, followed by synthesis, also known as the frame operator,
$S:\mathbb{R}^{n}$ $\rightarrow\mathbb{R}^{n}$
$x$ $\mapsto\sum_{i\in I}\langle x,\phi_{i}\rangle\cdot\phi_{i}.$
In matrix notation, $S=DC$. The frame operator is positive and self-adjoint, and if $\Phi$ is a frame, it is additionally invertible. Hence, one can write any $x\in\mathbb{R}^{n}$ as
$x=S^{-1}Sx=\sum_{i\in I}\langle x,\phi_{i}\rangle\cdot S^{-1}\phi_{i}.$ (45)
We call $\tilde{\Phi}=\left(S^{-1}\phi_{i}\right)_{i\in I}$ the canonical dual frame for $\Phi$. Denoting the synthesis operator associated with $\tilde{\Phi}$ by $\tilde{D}$, then Equation (45) is equivalent to
$\tilde{D}Cx=x.$ (46)
In other words, $\tilde{D}$ is a left-inverse of $C$ (given by the pseudo-inverse of $C$) [10]. This can be thought of as reconstructing $x$ from its frame coefficients using the canonical dual frame $\tilde{\Phi}$. If $\Phi$ is redundant ($m>n$), there are infinitely many different possibilities of constructing a left-inverse of $C$. All of them can be interpreted as the synthesis operator of a (non-canonical) dual frame. Using this machinery, we define a reconstruction operator for $C_{\alpha}$ analogously to (46). Note that unless $I_{x}^{\alpha}\neq I$ for all $x\in K$, there is not one reconstruction operator for all $x\in K$.
Definition 4.
We define the ReLU-synthesis operator associated with the collection $\Phi=(\phi_{i})_{i\in I}\subset\mathbb{R}^{n}$, the bias $\alpha\in\mathbb{R}^{m}$, and the index set $J\subseteq I$ by
$\begin{split}D^{\alpha}_{J}:\mathbb{R}^{m}&\rightarrow\mathbb{R}^{n}\\ (c_{i})_{i\in I}&\mapsto\sum_{i\in J}\left(c_{i}+\alpha_{i}\right)\cdot\phi_{i}.\end{split}$ (47)
Note that $\mathbb{R}^{m}$ is fixed as domain, the sum, however, runs over the index set $J$. When using the ReLU-synthesis operator associated with a dual frame of $\Phi_{J}$ we obtain a reconstruction formula in the spirit of (46).
Theorem 8 (ReLU-dual I).
Let $\Phi$ be $\alpha$-rectifying on $K$ for $\alpha\in\mathbb{R}^{m}$ and choose $x_{0}\in K$. Let $\widetilde{\Phi}_{I_{x_{0}}^{\alpha}}=(\tilde{\phi}_{i})_{i\in I_{x_{0}}^{\alpha}}$ be any dual frame of $\Phi_{I_{x_{0}}^{\alpha}}$, and $\tilde{D}^{\alpha}_{I_{x_{0}}^{\alpha}}$ the associated ReLU-synthesis operator. Then for all $x\in K$ such that $I_{x_{0}}^{\alpha}\subseteq I_{x}^{\alpha}$ it holds that
$\tilde{D}_{I_{x_{0}}^{\alpha}}^{\alpha}C_{\alpha}x=x.$ (48)
Proof.
For $x\in K$ with $I_{x_{0}}^{\alpha}\subseteq I_{x}^{\alpha}$, the operator composition in (48) reduces to the usual frame decomposition with $\Phi_{I_{x_{0}}^{\alpha}}$,
$\tilde{D}_{I_{x_{0}}^{\alpha}}^{\alpha}C_{\alpha}x$ $=\sum_{i\in I_{x_{0}}^{\alpha}}\left(\max(0,\langle x,\phi_{i}\rangle-\alpha_{i})+\alpha_{i}\right)\cdot\tilde{\phi}_{i}$
$=\sum_{i\in I_{x_{0}}^{\alpha}}\langle x,\phi_{i}\rangle\cdot\tilde{\phi}_{i}=x.$
Loosely speaking, the condition $I_{x_{0}}^{\alpha}\subseteq I_{x}^{\alpha}$ for a reference vector $x_{0}\in K$ means that we may use the left-inverse $\tilde{D}_{I_{x_{0}}^{\alpha}}^{\alpha}$ for all input elements $x$ that share at least all active elements with $\Phi_{I_{x_{0}}^{\alpha}}$. By re-writing the condition on the level of index sets, we can alternatively choose a reference sub-space via an index set $J$ instead of fixing a reference vector $x_{0}$.
Corollary 3 (ReLU-dual II).
Let $\Phi$ be $\alpha$-rectifying on $K$ for $\alpha\in\mathbb{R}^{m}$ and choose $J\subseteq I$ such that $\Phi_{J}$ is a frame. Let $\tilde{\Phi}_{J}$ be a dual frame and $\tilde{D}_{J}^{\alpha}$ the associated ReLU-synthesis operator. Then for all $x\in K$ with $J\subseteq I_{x}^{\alpha}$ it holds that
$\tilde{D}_{J}^{\alpha}C_{\alpha}x=x.$
The approach in the above corollary is particularly useful in the context of the bias estimation procedures described in Section 3.2: Given a decomposition of $\Phi$ into sub-frames, either by all different most correlated sub-frames (29) or via the facets of $P_{\Phi}$ (41), we can compute all associated left-inverses in advance and use them for reconstruction on demand.
The numerical implementation of a ReLU-dual is straightforward when using canonical duals of the sub-frames $\Phi_{J}$. Corresponding pseudo-code can be found in the appendix.
Excursion: Reconstruction from PReLU-layers
There are various modifications of the ReLU activation function. One of them is the parametrized ReLU, or PReLU, given by $\operatorname{PReLU}_{\gamma}=\max(\gamma s,s)$ with $0<\gamma\leq 1$ [17]. As this is an injective function, the associated PReLU-layer with weights given by $\Phi$ and any bias $\alpha$ is injective if and only if $\Phi$ forms a frame. In this case, we obtain a PReLU-dual by
$\begin{split}\tilde{D}^{\alpha}_{\gamma}:\mathbb{R}^{m}&\rightarrow\mathbb{R}^{n}\\ (c_{i})_{i\in I}&\mapsto\sum_{i\in I_{x}^{\alpha}}\left(c_{i}+\alpha_{i}\right)\cdot\tilde{\phi}_{i}+\sum_{i\notin I_{x}^{\alpha}}\gamma^{-1}\left(c_{i}+\alpha_{i}\right)\cdot\tilde{\phi}_{i},\end{split}$ (49)
where $\widetilde{\Phi}=(\tilde{\phi}_{i})_{i\in I}$ is any dual frame for $\Phi$.
4.2 Stability of Reconstruction
Now that we have a reconstruction mapping, we discuss how stable it is in a Lipschitz-sense. For this, we perform a stability analysis in analogy to how it was done for the phase retrieval operator in [9] and the saturation operator in [2]. Note that this is in contrast to the usual stability analysis for neural networks, where the Lischitz-stability of the forward mapping is considered.
Definition 5.
A frame $\Phi$ does $\kappa$-stable $\alpha$-rectification if there is $\kappa>0$ s.t.
$\|y-z\|^{2}\leq\kappa\cdot\|C_{\alpha}y-C_{\alpha}z\|^{2}$ (50)
holds for all $y,z\in K$.
If $\Phi$ is $\alpha$-rectifying then the frame inequality (4) for $C_{\alpha}$ holds immediately, i.e., there are constants $0<A_{\alpha}\leq B_{\alpha}<\infty$ such that
$A_{\alpha}\cdot\|x\|^{2}\leq\|C_{\alpha}x\|^{2}\leq B_{\alpha}\cdot\|x\|^{2}$ (51)
for all $x\in\mathbb{R}^{n}$. In [7] it was shown that the largest lower bound $A_{\alpha}$ for (51) coincides with the smallest eigenvalue among the frame operators of all possible active sub-frames $\Phi_{I_{x}^{\alpha}}$ with $x\in K$. Consequently $\alpha$-rectifying frame does $A_{\alpha}^{-1}$-stable $\alpha$-rectification on $K$. An analog statement holds for the upper bound $B_{\alpha}$.
As $A_{\alpha}$ might be very small the problem may become globally ill-conditioned. To get a better understanding of how stable the reconstruction process is for smaller portions of the data, we extend the discussion in [7] by proposing to consider local stability constants additionally.
Definition 6.
For $x_{0}\in K$, a frame $\Phi$ does $\kappa_{x_{0}}$-stable $\alpha$-rectification near $x_{0}$ if there is $\varepsilon>0$ and $\kappa_{x_{0}}=\kappa(x_{0})>0$ such that
$\|y-z\|^{2}\leq\kappa_{x_{0}}\cdot\|C_{\alpha}y-C_{\alpha}z\|^{2}$ (52)
holds for all $y,z\in\mathring{B}_{\varepsilon}(x_{0}).$
It turns out that, locally, we can in general only guarantee that an $\alpha$-rectifying frame does stable $\beta$-rectification for $\beta<\alpha$.
Proposition 5.
Let $\Phi$ be $\alpha$-rectifying on $K$. For $x_{0}\in K$ let $J=J(x_{0})\subseteq I_{x_{0}}^{\alpha}$ be such that $\Phi_{J}$ is a frame with lower frame bound $A_{J}$. Then $\Phi$ does $A_{J}^{-1}$-stable $\beta$-rectification near $x_{0}$ for $\beta<\alpha$.
Proof.
Let $x_{0}\in K$. There is $J=J(x_{0})\subseteq I_{x_{0}}^{\alpha}$ such that $\Phi_{J}$ is a frame with frame bounds $0<A_{J}\leq B_{J}<\infty$. Analog to the global case, the bi-Lipschitz condition
$A_{J}\cdot\|y-z\|^{2}\leq\|C_{\alpha}y-C_{\alpha}z\|^{2}\leq B_{J}\cdot\|y-z\|^{2}$ (53)
holds for all $y,z\in K$ with $I_{y}^{\alpha},I_{z}^{\alpha}\supseteq J$. Let $\beta\in\mathbb{R}^{m}$ such that $\langle x_{0},\phi_{i}\rangle\geq\alpha_{i}>\beta_{i}$ for all $i\in J$. Hence, there is $\varepsilon>0$ sufficiently small such for $y,z\in\mathring{B}_{\varepsilon}(x_{0})$ we have that $I_{y}^{\beta},I_{z}^{\beta}\supseteq J$. Since (53) still holds for $C_{\beta}$, we get that $\Phi$ does $A_{J}^{-1}$-stable $\beta$-rectification near $x_{0}$. ∎
Corollary 4.
Let $\Phi$ be $\alpha$-rectifying on $K$, and $x_{0}\in K$ such that $\langle x_{0},\phi_{i}\rangle>\alpha_{i}$ for all $i\in J$. Then $\Phi$ does $A_{J}^{-1}$-stable $\alpha$-rectification near $x_{0}$.
The computation of a lower frame bound of $\Phi_{J}$ is numerically very unstable, especially if $\Phi_{J}$ is contained in the collection of active frame elements. Under certain circumstances we can get a lower frame bound of $\Phi_{J}$ in terms of the bias values $\alpha_{i}$, $i\in J$ for free.
Lemma 7.
Let $\Phi$ be $\alpha$-rectifying on a bounded $K$ with $M=\sup_{x\in K}\|x\|$, and $J\subseteq I$ such that $\Phi_{J}$ is still a frame. If $\alpha_{i}\geq 0$ for all $i\in J$ and there is at least one $i_{0}\in J$ with $\alpha_{i_{0}}>0$ then $A_{J}^{\alpha}=M^{-2}\sum_{i\in J}\alpha_{i}^{2}$ is a lower frame bound for $\Phi_{J}$.
Proof.
The statement follows from $\sum_{i\in J}|\langle x,\phi_{i}\rangle|^{2}\geq M^{-2}\|x\|^{2}\sum_{i\in J}\alpha_{i}^{2}>0$. ∎
Hence, under the conditions above, $A_{I_{x}^{\alpha}}^{\alpha}$ yields a good indicator if the collection of frame elements that are active for $x$ and $\alpha$ is well conditioned.
4.3 Image of ReLU-layers
In the context of our work, it is specifically interesting to know what the image of a ReLU-layers looks like. The impact is two-folded.
  • (1)
    Sampling from the image of a ReLU-layer and reconstructing them yields a valuable method to generate new consistent data.
  • (2)
    To study the injectivity of a ReLU-layer that applies to the output of a previous one it is crucial to know where the image of the input domain $K$ lies under the ReLU-layer.
We give a partial answer to this question in the following. For bounded $K$ we can use the upper bound in (51) to find the smallest closed non-negative ball in $\mathbb{R}^{m}$ that contains $C_{\alpha}\left(K\right)$.
Lemma 8.
Let $\Phi$ be a frame and let $B_{0}$ denote the largest optimal upper frame bound among all sub-frames $\Phi_{I_{x}^{\alpha}}$ with $x\in K$ and let $M=\sup_{x\in K}\|x\|$. Then
$C_{\alpha}\left(K\right)\subseteq\mathbb{B}_{\sqrt{B_{0}}M}^{+},$ (54)
where $\sqrt{B_{0}}M$ is the minimal radius.
Proof.
By (51), for $x\in K$ we have that $\|C_{\alpha}x\|^{2}\leq B_{0}\|x\|^{2}\leq B_{0}M^{2}$. The application of the ReLU function then corresponds to the projection onto the non-negative part of $\mathbb{B}_{\sqrt{B_{0}}M}$. Since all estimations are sharp, the claim follows. ∎
As already mentioned, understanding the images of specific sets under ReLU-layers is crucial to understand how data is processed and passed on to a next layer. The above lemma is just a small step toward revealing this knowledge which can be used for unraveling certain behaviors of neural networks, and further, enhancing their interpretability and transparency.
5 Conclusion
This manuscript studies the injectivity of ReLU-layers and the exact recovery of input vectors from their output using a frame theoretic perspective. Among many basic properties and insights about ReLU-layers on bounded domains, the main theoretical results are three different characterizations of the injectivity of ReLU-layers that together provide a complete picture of its injectivity behavior as a non-linear deterministic operator. A significant portion of the research focuses on the computation of a maximal bias for a given frame $\Phi$ and a domain $K$, such that the associated ReLU-layer is injective on $K$. This characterization is particularly interesting as it allows for the theoretical results to be applied in practical applications easily. We discuss two different methods to approach this, both of which have distinct advantages and disadvantages, and provide algorithmic solutions to compute approximations of a maximal bias in practice.
The second part of this work is devoted to the derivation of reconstruction formulas for injective ReLU-layers, based on the concept of duality in frame theory. A stability analysis of the reconstruction operator completes the discussion.
In summary, this paper provides fundamental theoretical results on how to study the channeling of information in ReLU-layers for given input data, made possible by using frame theory as a tool. The results are intended to be designed in a general, yet, accessible way such that they may stimulate further theoretical research, but are also directly applicable in real-world settings. It is remarkable how diverse the approaches for studying this simple activation function can be, and we are happy to be able to contribute to a better understanding of these fundamental and ubiquitous building blocks of neural networks. Of course, there are many more aspects that are yet to be explored. In particular, we hope that our work will inspire people to tackle and, moreover, solve the main question related to the ReLU activation, namely how the information flow in ReLU-networks works, i.e., how can we characterize the injectivity of the concatenation of ReLU-layers.
Acknowledgement
D. Haider is recipient of a DOC Fellowship of the Austrian Academy of Sciences at the Acoustics Research Institute (A 26355). The work of P. Balazs was supported by the FWF projects LoFT (P 34624) and NoMASP (P 34922). The authors would particularly like to thank Daniel Freeman for his valuable input during very enjoyable discussions and Hannah Eckert for her work on the numerical experiments.
Appendix
A - Remarks on Admissibility and Directed Spanning Sets
With this short comment, we aim to complete the circle between three perspectives to characterize the injectivity of a ReLU-layer on $\mathbb{R}^{n}$.
In the work by Bruna et al. [7] the injectivity of a ReLU-layer (there called half-rectification operator) was linked to an admissibility condition of a $J\subseteq I$. There, $J$ is called admissible for $\Phi$ and $\alpha$ if
$\bigcap_{i\in J}\{x\in\mathbb{R}^{n}:\langle x,\phi_{i}\rangle>\alpha_{i}\}\cap\bigcap_{i\notin J}\{x\in\mathbb{R}^{n}:\langle x,\phi_{i}\rangle<\alpha_{i}\}\neq\emptyset.$
Puthawala et al. in [25] already pointed out that Proposition $2.2$ in [7] is not exactly related to the injectivity of a ReLU-layer. However, with a slight modification to
$\bigcap_{i\in J}\{x\in\mathbb{R}^{n}:\langle x,\phi_{i}\rangle\geq\alpha_{i}\}\cap\bigcap_{i\notin J}\{x\in\mathbb{R}^{n}:\langle x,\phi_{i}\rangle<\alpha_{i}\}\neq\emptyset,$
this stands in direct relation to the index sets $I_{x}^{\alpha}$, as introduced in Definition 2 in the present manuscript. Indeed, with this modified definition, $J$ is $\alpha$-admissible for $\Phi$ if and only if there is $x\in\mathbb{R}^{n}$ such that $J=I_{x}^{\alpha}$. As a consequence, the equivalence of $(i)$ and $(ii)$ in Corollary 1 in the present manuscript with $K=\mathbb{R}^{n}$, Proposition $2.2$ in [7] with the modified admissibility condition and Theorem $2$ in [25] are equivalent.
B - Omnidirectionality
For the sake of completing the comment about omnidirectionality in Part B. of Section 3.2 we formally prove that by adding a single vector, any non-omnidirectional frame can be made omnidirectional and how omnidirectionality can be checked numerically.
Lemma 9.
Let $\Phi=(\phi_{i})_{i\in I}$ be a non-omnidirectional frame, then
$\Phi^{\prime}=\left(\Phi,-\frac{\sum_{i\in I}\phi_{i}}{\|\sum_{i\in I}\phi_{i}\|}\right)$
is omnidirectional.
Proof.
At first, note that if $\sum_{i\in I}\phi_{i}=0$, then $\Phi$ is already omnidirectional. Let $c_{1},...,c_{m}=\frac{1}{1+\|\sum_{i\in I}\phi_{i}\|}$ and $c_{m+1}=\frac{\|\sum_{i\in I}\phi_{i}\|}{1+\|\sum_{i\in I}\phi_{i}\|}$ and $\phi_{m+1}=-\frac{\sum_{i\in I}\phi_{i}}{\|\sum_{i\in I}\phi_{i}\|}$, then
$\sum_{i=1}^{m+1}c_{i}\cdot\phi_{i}=0.$ (55)
Note that $c_{i}>0$ for all $i=1,\dots,m+1$ and $\sum_{i=1}^{m+1}c_{i}=1$. Hence, in (55) we wrote 00 as convex combination of all elements of $\Phi^{\prime}$. This implies that $0\in\mathring{P}_{\Phi^{\prime}}$, hence $\Phi^{\prime}$ is omnidirectional. ∎
Regarding the verification of omnidirectionality, let $D$ be the synthesis matrix associated with $\Phi$, i.e., it consists of the column vectors $\phi_{i}$ for $i\in I$. Then, verifying omnidirectionality is equivalent to the existence of a solution for the convex optimization problem $\min\ \|Dc\|\ \text{subject to}\ c>0$.
Figure 9: Left: The frame is omnidirectional. Right: The frame is not omnidirectional.
C - Algorithms
In this part we provide detailed pseudo-code for all algorithms in the paper to facilitate their implementation. Our Python implementations can be found under https://github.com/danedane-haider/Alpha-rectifying-frames.
C1. Sampling-based Bias Estimation. Given a frame $\Phi$, a data domain $K$, and a sequence of samples $X_{N}\subset K$, Algorithm 1 demonstrates the sampling-based bias estimation presented in Theorem 6. The samples $x_{k}\in X_{N}$ can be random, e.g., $x_{k}\sim\mathcal{U}(K)$ for suitable $K$. Assuming the frame to be full-spark, then $J^{*}(x_{k})$ consists of the indices for the largest $n$ frame coefficients $(\langle x_{k},\phi_{i}\rangle)_{i\in I}$.
Algorithm 1 Sampling-based approach for approximating $\alpha_{K}^{\flat}$
initialize $(\alpha^{(0)})=\alpha_{\Phi}^{\Delta}$, $k=0$
for $z$ in $X_{N}$ do
     compute $(\langle z,\phi_{i}\rangle)_{i\in I}$
     get $J^{*}(z)$
     update $(\alpha^{(k+1)})_{i}\leftarrow\min\{\langle z,\phi_{i}\rangle,(\alpha^{(k)})_{i}\}$ for all $i\in J^{*}(z)$
     $k=k+1$
end for
The for-loop can be replaced with a while-loop conditioned on $\|\alpha^{(k+steps)}-\alpha^{(k)}\|>\varepsilon>0$, where the parameter $steps$ determines how many updates should be done before checking the condition.
C2. Polytope Bias Estimation. Given a frame $\Phi$ and a bounded domain $K$. The vertex-facet incidences $I_{F_{j}}$ for $P_{\Phi}$ can computed with convex hull algorithms, e.g., using the property simplices from scipy.spatial.ConvexHull in Python. In the following we demonstrate how to compute the biases from Proposition 4.
$(i)$ $K=\partial P_{\Phi}:$ Recall that $\alpha_{\Phi}^{\Delta}$ is given by
$\left(\alpha_{\Phi}^{\Delta}\right)_{i}=\min_{\begin{subarray}{c}\ell\in I_{F_{j}}\\ j:\phi_{i}\in F_{j}\end{subarray}}\langle\phi_{\ell},\phi_{i}\rangle.$ (56)
Algorithm 2 PBE for $\partial P_{\Phi}$
compute $I_{F_{j}}$ for all facets
for $j=1,\dots,\#$facets do
     $\beta_{j}=\min_{k<\ell\in I_{F_{j}}}\langle\phi_{k},\phi_{\ell}\rangle$
end for
for $i=1,\dots,m$ do
     $\left(\alpha_{\Phi}^{\Delta}\right)_{i}=\min\limits_{j:i\in I_{F_{j}}}\beta_{j}$
end for
$(ii)$ $K=\mathbb{S}:$ Recall that $\alpha_{\mathbb{S}}^{\Delta}$ is given by
$\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i}=\min\{\min_{\begin{subarray}{c}y\in F_{j}^{\mathbb{S}}\\ j:\phi_{i}\in F_{j}\end{subarray}}\langle y,\phi_{i}\rangle,\left(\alpha_{\Phi}^{\Delta}\right)_{i}\}.$ (57)
First, we show that for fixed $i\in I_{F_{j}}$,
$\min_{\begin{subarray}{c}y\in F_{j}^{\mathbb{S}}\\ j:\phi_{i}\in F_{j}\end{subarray}}\langle y,\phi_{i}\rangle$ (58)
can be computed using convex linear programs. Letting $D_{I_{F_{j}}}$ denote the synthesis operator associated with $\Phi_{I_{F_{j}}}$ then we can write any $x\in F_{j}^{\mathbb{S}}$ as $x=\sum_{\ell\in I_{F_{j}}}c_{\ell}\phi_{\ell}=D_{I_{F_{j}}}c.$ So for any $i\in I_{F_{j}}$ the solution of (58) is found by solving the linear program
$\begin{split}\min\ \left(y^{\top}D_{I_{F_{j}}}\right)c\\ \text{subject to}\ c&\geq 0\\ \|D_{I_{F_{j}}}c\|_{2}&=1.\end{split}$ (59)
If $\left(\alpha_{\Phi}^{\Delta}\right)_{i}<0$, then the above minimum is negative since $\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i}\leq\left(\alpha_{\Phi}^{\Delta}\right)_{i}<0$. Hence, we can replace $\|D_{I_{F_{j}}}c\|_{2}=1$ by $\|D_{I_{F_{j}}}c\|_{2}\leq 1$ making the problem convex.
Algorithm 3 PBE for $\mathbb{S}$
compute $\alpha_{\Phi}^{\Delta}$
for $i=1,\dots,m$ do
     if $\left(\alpha_{\Phi}^{\Delta}\right)_{i}\geq 0$ then
         $\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i}\leftarrow\left(\alpha_{\Phi}^{\Delta}\right)_{i}$
     else
         $\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i}\leftarrow$ solution of (59)
     end if
end for
$(iii)$ $K=\mathbb{D}_{r,s}:$ Let $0\leq s<r$ and recall that $\alpha_{\mathbb{B}}^{\Delta}$ is given by
$\left(\alpha_{\mathbb{B}}^{\Delta}\right)_{i}=\min\{s,\left(\alpha_{\mathbb{S}}^{\Delta}\right)_{i}\}.$ (60)
One gets the general case by scaling with $r^{-1}$. The case $s=0$ yields a bias estimation for $\mathbb{B}_{r}$. Since (60) depends on $\alpha^{\mathbb{S}}$ in a trivial way, we omit the algorithm.
$(iv)$ $K=\mathbb{B}^{+}_{r}:$ Let $e\in\mathbb{R}^{m}$ be arbitrary and recall that $\alpha^{\mathbb{B}^{+}}$ is given by
$\alpha_{i}^{\mathbb{B}^{+}}=\begin{cases}\alpha_{i}^{\mathbb{B}}&\text{ for }i\in I^{+}\\ e_{i}&\text{ else,}\end{cases}$ (61)
The crux here is to compute the index set $J^{+}=\{j\in I:F_{j}\cap\mathbb{R}^{n}_{+}\neq\emptyset\}$ which leads to $I^{+}=\bigcup_{j\in J^{+}}I_{F_{j}}$. One way to verify $F_{j}\cap\mathbb{R}^{n}_{+}\neq\emptyset$ is to check the feasibility of the convex optimization problem
$\min\ \|D_{I_{F_{j}}}c\|_{2}$
subject $\text{subject to}\ c$ $\geq 0$ (62)
$\sum_{i}c_{i}$ $=1.$
If (C - Algorithms) has a solution, then there is $c\in\mathbb{R}^{n}_{+}$ that can be written as a convex linear combination of the vertices of $F_{j}$, hence, $F_{j}\cap\mathbb{R}^{n}_{+}\neq\emptyset$.
C3. ReLU-duals and Reconstruction. For $J\subseteq I$ we denote by $C_{J}$ the analysis operator for the collection $\Phi_{J}$. This corresponds to the $|J|\times n$ matrix $C_{J}$ consisting of the row vectors $\phi_{i}$ for $i\in J$. Algorithm 4 describes how to build the synthesis matrices to construct the ReLU-duals from Corollary 3.
Algorithm 4 Construction of the matrix for ReLU-duals
choose $J\subseteq I$ such that $\Psi=\Phi_{J}$ is a frame
$S^{-1}_{J}\leftarrow\left((C_{J})^{\top}C_{J}\right)^{-1}$
$\tilde{D}_{J}\leftarrow\begin{pmatrix}|&|&&|\\ \\ S^{-1}_{J}\psi_{1}&S^{-1}_{J}\psi_{2}&\cdots&S^{-1}_{J}\psi_{|J|}\\ \\ |&|&&|\end{pmatrix}$
Note that $\tilde{D}_{J}$ is the synthesis operator associated with a non-canonical dual frame for $\Phi$ in the classical sense.
Given $z=C_{\alpha}x_{0}$ we can read off $I_{x_{0}}^{\alpha}$ from $z$ directly. Assuming that we have a suitable list of index sets of sub-frames (e.g., all most correlated bases of all facet sub-frames), finding those that are contained in $I_{x_{0}}^{\alpha}$ is easy. Choose one, say $J$. The choice $J=I_{x_{0}}^{\alpha}$ is valid too. Then the matrix $\tilde{D}_{J}$ provides reconstruction as following.
Algorithm 5 Applying the ReLU-dual: Reconstruction of $x_{0}$ from $z=C_{\alpha}x_{0}$
choose $J\subseteq I_{x_{0}}^{\alpha}$ such that $\Phi_{J}$ is a frame
unbias $z\leftarrow z+\alpha$
restrict $\zeta\leftarrow(z_{i})_{i\in J}$
reconstruct $\tilde{D}_{J}\zeta=x_{0}$